### v1

# Galaxy formation with Wave/Fuzzy Dark Matter: The core-halo structure

Abstract
Dark matter-dominated cores have long been claimed for the well-studied local group dwarf galaxies. More recently, extended stellar halos have been uncovered around several of these dwarfs through deeper imaging and spectroscopy. Such core-halo structures are not a feature of conventional cold dark matter (CDM), based on collisionless particles where smooth, scale-free profiles are predicted. In contrast, smooth and prominent dark matter cores are predicted for Warm and Fuzzy/Wave Dark Matter (WDM/DM) respectively. The question arises to what extent the visible stellar profiles should reflect this dark matter core structure. Here we compare cosmological hydrodynamical simulations of CDM, WDM DM, aiming to predict the stellar profiles for these three DM scenarios. We show that cores surrounded by extended halos are distinguishable for WDM and DM, with the most prominent cores in the case of DM, where the stellar density is enhanced in the core due to the presence of the relatively dense soliton. Our analysis demonstrates that such behavior does not appear in CDM, implying that the small-scale cut-off in the power spectrum present for WDM and DM provides a core-halo transition. Consequently, we estimate the mass of the DM particle at this core-halo transition point. Furthermore, we observe the anticipated asymmetry for DM due to the soliton’s random walk, a distinctive characteristic not found in the symmetric distributions of stars in Warm and CDM models.
keywords:
Dark Matter, Galaxies: dwarfs, Galaxies: kinematics and dynamics1 Introduction
Dark Matter is understood to be non-relativistic, i.e. cold, as required to form galaxies gravitationally and for explaining the spectrum of CMB fluctuations in detail (Collaboration et al., 2018). However, the standard heavy particle interpretation for Cold Dark Matter (CDM) faces a stringent laboratory absence of any new particle signature (Aprile et al., 2018; Collaboration, 2020) and several inconsistencies have emerged between CDM predictions and the properties of dwarf galaxies (Moore, 1994; de Blok, 2010; Marsh & Pop, 2015; Klypin et al., 1999; Safarzadeh & Loeb, 2021). Warm Dark Matter (WDM) is defined to be relativistic initially with a keV-scale mass that is sufficiently redshifted to become non-relativistic before galaxy formation, generating smooth cores and suppressing formation of low-mass galaxies. Alternatively, for very light bosons of , an inherently non-relativistic possibility of dark matter as a Bose-Einstein condensation (Widrow & Kaiser, 1993; Hu et al., 2000), originally termed Fuzzy Dark Matter (FDM), has been emerged by the first simulations in this context, which predict prominent solitonic standing wave cores surrounded by pervasive interference on the de Broglie wavelength scale within all galaxies and filaments (Schive et al., 2014a; Mocz et al., 2017; Veltmaat et al., 2018; Hui, 2020). This was then termed as ‘Wave Dark Matter’, DM, (Schive et al., 2014a; Hui et al., 2021). While DM predicts almost identical large-scale structure to CDM in terms of the network of filaments and clusters (Schive et al., 2014a; Mocz et al., 2017), the small-scale structure of DM halos is very different in DM paradigm, as DM cannot be confined to less than the de Broglie scale. Consequently, in DM, with a light boson mass of , the dwarf galaxy formation is suppressed and there is also a prominent solitonic core (Schive et al., 2014a, b; Mocz et al., 2017; Veltmaat et al., 2018; Niemeyer, 2020) in the galaxy.
On these scales, self-gravity balances the effective pressure from the Uncertainty Principle in the ground state at the de Broglie wavelength. Crucially, smaller galaxies are predicted to have wider cores of lower density because the soliton is larger at lower momentum, which we test here.
Many independent claims of dark matter and stellar cores in deep images and spectroscopy are now available for the well-studied, Local Group dwarf spheroidal galaxies, dSph, providing one of the main sources of tension for CDM for which the lack of any predicted cores reflects the scale-free nature of simplest, collisionless DM scenario. More recently the presence of extended stellar halos has been uncovered in several local dwarfs, independently, with deep imaging and spectroscopy (Chiti et al., 2021; Collins et al., 2021). A joint analysis of the available data now suggests that such a core-halo structure with a sharp density transition between the core and the halo is a common feature of both dSph galaxies and also the smaller new class of UFG (Ultra faint galaxies) (Pozo et al., 2020, 2023). This core-halo behaviour motivates the examination of hydrodynamical simulations with a comparison of the differing DM contenders as a possible distinguishing feature.
For cold dark matter (CDM), the scale-free formation of DM structures is predicted to extend to arbitrarily small scales, reaching lower halo masses with higher DM concentrations. In contrast, small-scale suppression of the power spectrum is inherent to both DM and WDM. In the case of DM, this suppression limits the minimum scale of structure to below the de Broglie length, which is determined by the boson mass. Consequently, DM structures below are suppressed for a choice of boson masses of eV, motivated by the observed kpc scale of dwarf galaxy cores. This limitation has a direct consequence for the formation timescales between CDM and DM, resulting in a delay in galaxy formation for DM compared to CDM. In both cases, we observe that the formation of the first "filamentary" structures occurs earlier in CDM, where they are comprised of low-mass subhaloes. In contrast, in DM/WDM, there is no fragmentation along the filaments due to the small-scale cutoff of the power spectrum. Instead, dark matter is distributed more continuously along these filaments.
The main difference between warm dark matter (WDM) and DM is that WDM exhibits a clear caustic structure in the distribution of dark matter, while DM is characterized by wave interference patterns that manifest as fringes in simulations. These fringes are now understood to generate turbulent density structures within halos, resulting from the superposition of numerous plane waves encoding the velocity dispersion in the halo (Mocz et al., 2020). In DM, these filaments persist for longer periods, as the gravitational attraction from the early formation of the first halos in cold dark matter (CDM) tends to disrupt them earlier. Moreover, the prolonged lifetimes of these filaments in wave dark matter favor the formation of a greater number of stars within them, leading to significant differences in the location and extent of stellar profiles in CDM and DM galaxies. This also causes baryonic objects to appear more diffuse or smoothed compared to CDM (Mocz et al., 2019, 2020), indicating that they could serve as excellent tracers of dark matter in DM. On the scale of solitons, the quantum pressure in DM can become sufficiently strong to counteract the self-gravity of the dark matter. This results in the formation of prominent solitonic standing waves at the base of each virialized potential, characterized by a spherical soliton core of a few kiloparsecs at the center of every dark matter halo (Schive et al., 2014a), in contrast to the much denser cusps observed in CDM (Navarro & Frenk, 1996) and WDM (Mocz et al., 2019).
In dSph galaxies it is known that stellar motions are dominated by the dark matter gravitational potential and so we expect some visible differences for stars in these galaxies depending on the nature of dark matter. In addition, the formation and early evolution of galaxies and filaments will also be sensitive to the nature of DM, arising in particular from the suppressed form of the high-k power spectrum for the WDM/DM models. The wave behaviour and the quantum pressure of DM provide different virialized DM halo structures and prominent soliton cores, unlike the smooth cores of WDM. While for CDM and WDM cuspy profiles are expected (Navarro & Frenk, 1996; Lovell et al., 2014), DM simulations of merging DM haloes clearly show the formation of soliton structures at the de Broglie wavelength scale (Schive et al., 2014a; Mocz et al., 2017; Schwabe et al., 2016), and are prominent in terms of the core density, standing above the surrounding DM halo. This is in contrast to the early fragmentation of filaments and cuspy haloes in CDM (Mocz et al., 2017, 2019, 2020). Mocz et al. (2019, 2020) demonstrated that baryonic feedback does not have a significant effect in halos of to for redshifts , in the sense of being unable to soften the cuspy profiles of CDM/WDM to produce cores. Consequently, neither CDM/WDM are able to explain the origin of the claimed cores from the dynamical studies of dSph galaxies (Read & Gilmore, 2005; Amorisco et al., 2013).
In this paper we compare the galaxy formation process between CDM, WDM, and DM, in the high-resolution cosmological simulations by Mocz et al. (2019, 2020, 2017). Here, we will consider the distribution of stars modelled in these simulations, where in the WDM and DM contexts these simulations have already reported stars to be born along dense dark matter filaments, tracing dark matter, and highlighted as “a smoking gun” signature of DM (Mocz et al., 2019). Here we will further examine these lowest redshifts, focusing on the question of core-halo structure for comparison with new deep imaging and spectroscopic data for the Milky Way and other local group dwarfs. For these objects, the data now indicates non-smooth profiles in these dark matter-dominated dwarfs as this stellar structure may potentially have important implications for the nature of dark matter. Moreover, we will look into the different evolution of the stellar populations and the impact of the three DM models on their spatial distribution.
The structure of this paper is as follows: firstly, in sections 2 and 3, we describe the radial structure and internal dynamics of the DM halo for comparison with the data. Next, in section 4, we present the simulation setup. In section 5, we focus on presenting all the results of the work, with particular emphasis on the main finding related to the detection of the core-halo structure. Finally, in section 6, we discuss our conclusions regarding the non-detection of the core-halo structure for CDM and its implications.
2 The Wave Dark Matter Halo
Ultralight bosons, such as Axions, have been considered as an ingredient in describing the DM (Widrow & Kaiser, 1993; Hu et al., 2000). In the absence of any self-interactions, the boson mass is the only free parameter that describes the DM. For a sufficiently light boson mass, the de-Broglie wavelength exceeds the mean free path leading the Axion to satisfy the ground state condition for a Bose-Einstein condensation. In the comoving coordinate, the governing equation for Axion reads as:
where is the wave function, is the gravitation potential and (1/(1+z)) is the cosmological scale factor. The system is normalized to the time scale , and to the scale length , with , where is the current density parameter (Widrow & Kaiser, 1993).
Recently, it has proved possible with advanced GPU computing to make reliable, high dynamic range cosmological simulations that solve the above equations, (Schive et al., 2014a; Schwabe et al., 2016; Mocz et al., 2017; May & Springel, 2021). These evolve to produce large-scale structures indistinguishable from CDM, but with virialized halos characterized by a solitonic core in the ground state that naturally explains the dark matter-dominated cores of dSph galaxies (Schive et al., 2014b). Another important feature arising from the simulations is that the central soliton is surrounded by an extended halo with a “granular” texture on the de-Broglie scale, due to interference of excited states, but which when azimuthally averaged follows the Navarro-Frenk-White (NFW) density profile (Navarro & Frenk, 1996; Woo & Chiueh, 2009; Schive et al., 2014a, b).
The fitting formula for the density profile of the solitonic core in a DM halo is obtained from cosmological simulations (Schive et al., 2014a, b) as:
where is the boson mass and is the solitonic core radius. The latter scales with the product of the galaxy mass and boson mass, obeying the following scaling relation which has been derived from simulations (Schive et al., 2014b):
Beyond the soliton, at radii larger than a transition radius (), the simulations also reveal that the halo roughly looks like NFW in shape, presumably reflecting the non-relativistic nature of condensates beyond the de Broglie scale, and therefore the total density profile can be written as:
In detail, the scale radius of the solitonic solution, which represents the ground state of the Schrodinger-Poisson equation, is related to the size of the halo through the uncertainty principle. From cosmological simulations, the latter is found to hold non-locally, relating a local property with a global one (for more details we refer to Schive et al. (2014b)).
3 Dynamical Model of Galaxies in Wave Dark Matter Halo
Classical dSph galaxies are known to be dominated by DM with stars being treated as the tracer particles (Gregory et al., 2019; McConnachie & Irwin, 2006; McConnachie et al., 2006; Kang & Ricotti, 2019) moving along the gravitational potential generated by the DM halo.
In this context, the corresponding velocity dispersion profile can be estimated by solving the spherically symmetric Jeans equation:
where is the DM halo mass obtained by integrating the spherically symmetric density profile in Eq. (5), is the anisotropy parameter (see (Binney & Tremaine, 2008), Equation (4.61)), and is the stellar density profile as given by:
where
Here, is the 3D scale radius of the stellar halo corresponding to the central stellar density, is the normalization of at the transition radius and the transition radius, , is the point where the solitonic structure ends and the halo begins at the juncture of the core and halo profiles.
4 Simulation setup
Throughout this manuscript, we extensively make use of the most recent simulations of DM of Mocz et al. (2019, 2020). These simulations adopt a simulation box of and a boson mass of and starts from an initial redshift of evolving to a final redshift of (5.56 for DM). The initial conditions are generated using the publicly available Boltzmann code AXION-CAMB. The selection of the final redshift is driven by the need for a resolution to achieve fully converged results. The DM spatial resolution is set at , accompanied by a baryonic resolution of . The cosmological parameters correspond to the Planck results.
The primary objective is to conduct a thorough comparative analysis of the stellar properties between the DM simulations and the ones stemming from both CDM and WDM simulations, spanning an extended duration down to . In the case of the WDM simulations, a particle resolution of was utilized, employing the identical hydrodynamical setup as the DM simulations. It is noteworthy that while the CDM simulation lacks any form of cutoff in its power spectrum, both the DM and WDM simulations adopt an initial power spectrum characterized by exponential suppression.
Baryons (gas and star formation) are included in the simulation, and couple to the dark matter gravitationally. The baryonic material experiences primordial and metal-line cooling, chemical enrichment, stochastic star formation with a density threshold of , supernova feedback via kinetic winds, and instantaneous uniform reionization at (Vogelsberger et al., 2013; Torrey et al., 2014; Pillepich et al., 2017; Springel et al., 2018), as has been used in the Illustris and Illustris-TNG projects. The stellar feedback model has previously been tuned to match key observables with CDM simulations (Vogelsberger et al., 2013). Whether FDM requires significant re-tuning of efficiency factors is left for future study. The subgrid model for feedback is meant to describe the effects of Type II supernovae (SNII), and uses the local star formation rate to set the mass loading of stellar winds driven by the energy available through SNII (Vogelsberger et al., 2013; Pillepich et al., 2018). The simulation also keeps track of metallicty, as described in Torrey et al. (2019). Star particles (representing a population of stars) are born with the metallicity of their surrounding interstellar gas. As star particles evolve in time, mass and metals from the aging stellar populations are returned to the interstellar medium, taking into account SN1a and SNII supernovae and AGB stars. Deposited metals in the gas evolve through passive advection with the fluid flow.
5 results
5.1 First Illustration of density profiles
Before delving into the intricacies of comparing various simulation types, we provide an initial depiction of the projected density profiles for CDM, WDM, and DM at two different redshifts. In Figure 1, we present the projected density profiles for both DM and stars. The first two rows illustrate CDM profiles, while the last two rows depict the corresponding profiles for WDM/DM. In each row, the left(right) panel presents (). Mocz et al. (2019) showed that WDM and DM exhibit remarkably similar evolutionary patterns across various redshifts, implying a resemblance in the development of large-scale structures. This stands in contrast to CDM. Consequently, in Figure 1 we exclusively focus on presenting the WDM results.
Figure 1 suggests slight discrepancies between galaxies in these three distinct DM scenarios. First, in the context of CDM, filaments rapidly collapse into compact halos, whereas in the case of DM, these filaments retain a smoother configuration without giving rise to small galaxies. Second, in the CDM scenario, stars seem to exhibit greater concentration with irregular structures. Conversely, in the case of DM and WDM, a flatter shape is evident, accompanied by galaxies possessing extended stellar halos. This imparts a visual impression that stars are positioned farther away from the center. As we consider individual stellar profiles of galaxies across distinct DM scenarios, a notable finding emerges. In the DM scenario, the stellar profiles exhibit a flatter behaviour in the inner regions, in alignment with the expected pattern of stars tracing the presence of a soliton in the core. In contrast, both WDM and CDM scenarios lack such a soliton, resulting in a cuspy inner profile for these scenarios. While the primary focus of this study lies in the comparison of evolved cosmological structures between DM and CDM, it is reasonable to anticipate that the large-scale structures in WDM would bear similarities to DM. As illustrated in Mocz et al. (2019), the resulting structures are remarkably akin for WDM and DM at z=5.56, despite the more pronounced differences in their internal halo structures.
5.2 Core-halo structure
Below, we explore the presence of core-halo structures across various galaxies.
5.2.1 Signature of the core-halo structure
We start off with comparing the stellar profiles at . Figure 2 compares the stellar density profiles of the three main galaxies in the DM model at with those of the CDM model. Moving from left to right, we present the first, second, and third galaxy, respectively, as was previously pointed out in Figure 1. In the first row, a notable difference is seen between the stellar profiles in the DM scenario and CDM. This highlights how stellar profiles can vary and evolve differently in distinct DM scenarios, suggesting they could serve as a valuable tool for analyzing DM signatures in dwarf galaxies. The second row provides a closer look at the inner part of the galaxy and offers a clear test to assess the compatibility of the profiles with a core-halo structure. To this end, we plot the ideal solitonic and NFW profiles that would correspond to the inner core and the characteristic asymptotic profile, respectively. It is worth noting that only in the case of DM do the NFW and solitonic profiles intersect while still fitting the stellar profile. This is never a possibility for any CDM galaxy. This result may indicate the potential presence of a primordial core-halo stellar structure. There is also a noticeable difference in the inner profile for galaxies two and three, which aligns with the presence or absence of a soliton, resulting in a core-cuspy disagreement. Furthermore, in the case of galaxy three, where both DM and CDM exhibit a similar halo with the same slope, we can observe how the NFW profile effectively fits the entire CDM stellar profile. This is a consequence of the anticipated cusp, whereas a soliton is required for DM to establish its flat inner profile. The reason why both stellar profiles appear to coincide with a flat core for galaxy one in the inner part could be attributed to the widely suggested notion that baryonic feedback may lead to the development of a core, even in the CDM scenario (Chan et al., 2015; Pontzen & Governato, 2012). This is more plausible at this redshift for G1 than for G2 and G3, given its greater mass as the inner slope of the DM halo profiles shows a strong mass dependence (Chan et al., 2015). It is likely the same reason why the core-halo structure had sufficient time to form only in this galaxy. The vertical line in the figure represents the comoving softening length, which is 0.19 kpc in this case. Below this radius, cautious must be taken in interpreting the data.
Next, we proceed with comparing the stellar profiles at between the WDM and CDM simulations, following the approach outlined in the previous section. It is worth noting that there was no available data at this redshift for DM. Despite the significant differences between WDM and DM on small scales where the core-halo structure is anticipated, their similar large-scale evolution and initial conditions could make the comparison between WDM and CDM valuable for future studies. The resulting extracted stellar profiles, shown in Figure 3, seem to reaffirm the same findings as the previous redshift for CDM. In CDM, there is no evidence or imprint of a core-halo structure identified in the halos of the three simulated galaxies. In contrast, one WDM halo exhibits a potentially compatible core-halo stellar profile, particularly in a primordial context, suggesting that they may still be in the process of forming for galaxies one and two. Furthermore, a well-defined core-halo structure is evident for the third galaxy, characterized by a notable density drop from the core to the halo, along with the distinctive change of regime point (transition point, ) marking the boundary between the inner core and the outer halo (as depicted in the lower panels of Figure 3). This is observable where the pure NFW and solitonic profiles intersect. It is interesting to observe that for WDM, the core-halo structure emerges earlier in a lower mass galaxy, such as galaxy three, in comparison to the more massive galaxy one. This structure should be established by a more pronounced core, likely a combination of a dense soliton and potentially baryonic feedback, both more prominent in heavier galaxies. Nevertheless, it is crucial to remember that WDM lacks a solitonic core. Indeed, the presence of a core-halo structure remains a topic for future analysis.
5.2.2 Stellar, dark matter and gas distribution
Here we compare distinct distributions of stellar, gaseous, and dark matter profiles in all three different simulated galaxies across different DM scenarios. The aim is to extend the former comparison made only for the stellar profiles, done in Figure 2 and 3, to further include the DM and gaseous distributions as well. Figure 4 presents these profiles at . Once again, the stark contrast between the stellar and DM profiles in each case is evident, with gas profiles exhibiting a more consistent pattern across different DM models. Notably, a common feature is the presence of a core-halo transition in the DM profile from DM and gas profiles, even if it is only visibly prominent in the case of stars in G1. In comparison with DM, WDM and CDM exhibit strikingly similar cuspy stellar and DM profiles, while the gas profile resembles that of the DM case. From the plot, it is inferred that in DM model both of the stellar and gaseous profiles demonstrate similar core radii and transition points compared with the DM profile. This suggests that stars, rather than gas, may hold the key to determine the underlying DM model. Another crucial observation is the nearly identical core size and transition point in stars and gas, closely mirroring that reproduced by the DM particles. This implies a potential baryonic-dark matter connection and hints at the possibility of baryonic tracking of dark matter particles (Pozo et al., 2021; Pozo et al., 2022; Pozo et al., 2020, 2023).
| Combinations | ||||
| (kpc) | (kpc) | eV | ||
| DM | 5.56 | |||
| WDM | 2.23 | |||
| dSph | 0 | |||
| dSph | 0 | |||
| dSph | 0 | |||
| Draco | 0 | - | ||
| Leo II | 0 | - | ||
| Sculptor | 0 | - | ||
| And IX | 0 | - | ||
| And XV | 0 | - |
| Halos | ||||
| kpc | kpc | |||
| DM | 8.3 | - | 52 | - |
| DM | 5.5 | - | 45 | - |
| DM | 5.9 | - | 46 | - |
| CDM | 17 | 39.5 | 65 | 83 |
| CDM | 4.5 | 26.8 | 42 | 73 |
| CDM | 8.3 | 26.8 | 52 | 73 |
| WDM | 13 | 35.5 | 60 | 80 |
| WDM | 7.3 | 38.2 | 49 | 82 |
| WDM | 6.9 | 24.8 | 48 | 70 |
5.2.3 Estimation of the core-halo parameters
We now proceed with MCMC-based DM profile fits (Eq. (7)) to the DM, as shown in Figure 5 ( Corner plot Figure 14). The core radius (), as defined in Eq. (4), plays a crucial role in characterizing the shape of the solitonic core profile. For the halo, we utilize a NFW profile with a scale radius () and the normalization factor (). The transition radius (), defining the point of density transition between the soliton and NFW profiles, is the only other free parameter. We vary this within a prior range determined by DM simulations, specifically to times (Schive et al., 2014a, b). The boson mass stands as the sole fixed parameter, maintaining a value of eV, consistent with that used for the simulations (Mocz et al., 2020). The best-fit yields a resulting profile with a total mass of , aligning well with the computed mass from Mocz et al. (2020) of for this galaxy at this redshift. In Figure 5 the blue dashed line presents the resulting profile when we conduct the analysis with the fixed total mass of , which corresponds to the total mass of that specific simulated halo (indicated by a red circle in Figure 1) at this redshift. It is worth noting that while it may not match the accuracy of the green profile with a freely varying mass, it still demonstrates excellent agreement. Furthermore, the resulting values for kpc and kpc, fall within the expected range for such dSphs, and are very close to the directly inferred soliton radius from the simulation ( kpc (Pozo et al., 2020, 2023)). This concordance between the suggested soliton size based on stellar behaviour and the actual size of the simulated DM halo soliton may hold significant implications for our understanding of how stellar behaviour is influenced by dark matter.
Next, we proceed with fitting a stellar core-halo profile for another case where a core-halo shape is evident: the WDM G3 at . It is crucial to note that in principle fitting a core-halo profile to WDM, without a physical soliton, might be challenging. In this sense, we are careful not to directly link the core-halo structure to a soliton but rather to an effective profile with a transition from the core to the halo in WDM at this lower redshift. The resulting values for core radius () as well as the transition point () are summarized in Table 1, alongside other pertinent variables, with the resulting profiles depicted in Figure 6 ( Corner plot Figure 15). It is noteworthy that these parameter values fall within the expected range when compared to those extracted from observed profiles of Local Group classical dSphs. This is particularly striking considering they were analyzed at different redshifts and are still in the process of the formation. Table 1 also provides the parameter values for the two extracted galaxies and the observed dSphs in the Local Group (Pozo et al., 2023). The core radius values are in excellent agreement with the mean values obtained for such dSphs in the Local Group.
The isolated galaxy (G1) from the DM simulation data (, Table 1) exhibits a core radius akin to that of two Milky Way satellite galaxies, Leo II and Sculptor. Conversely, the third galaxy from the WDM simulation data (WDM, Table 1) aligns exceptionally well with the observed cores of Sculptor and And IX. Nonetheless, there are slight disparities in the transition point, which is reasonable given the difference in redshift and the fact that these simulated galaxies are still in the process of formation. The observed increase in between and WDM could be attributed to the lower redshift of WDM, which aligns with the expectation of a larger . Notably, despite the significant redshift difference (almost 6 between the observed galaxies and the observable Local Group dwarfs), the values remain similar. This suggests that for a DM origin, the soliton is likely the first structure to form in this core-halo structure.
We have included the virial masses of all these three simulated galaxies at both redshifts in Table 2 to facilitate a clearer comparison of their evolution across different DM models. Additionally, this table serves as a useful tool for comparing the masses of G1 (isolated) with those in G2 and G3 within the filamentary structure, as depicted in Figure 1.
5.2.4 Comparison with observations
We proceed by comparing the extracted profile with two known dwarf galaxies, Cetus and Andromeda XXI, that orbit the Milky Way and Andromeda, respectively, and exhibit the most similar density gaps (See Figure 7). In the left panel, the profiles are normalized based on their stellar peak density, while the right panel is additionally normalized by their respective transition radii. The black profiles represent the mean stellar profile of all Local Group dwarf spheroidal galaxies (Pozo et al., 2023). The right panel demonstrates that the extracted DM stellar profile (blue profile) aligns well with a real observed stellar profile, Andromeda XXI (red profile). This indicates that the extracted profile can reproduce observed dSph structures, in contrast to the CDM profiles.
Furthermore, this panel reveals that the inner portions of the profiles are in good agreement, suggesting that they are consistent with a solitonic core. However, it’s crucial to note that Andromeda XXI and the extracted profile are at different redshifts, 5.5 and 0 respectively. This indicates that the extracted profile still has much more time to evolve, and its final core-halo structure at redshift 0 could substantially change over time.
| Halos | (-2<[Fe/H]<0) | (-4<[Fe/H]<-2) | (-6<[Fe/H]<-4) | (-8<[Fe/H]<-6) | |
| DM | 12340 | 93.12 | 4.91 | 0.71 | 1.26 |
| DM | 1950 | 80.62 | 14.82 | 1.44 | 3.12 |
| DM | 1550 | 77.22 | 16.91 | 1.68 | 4.19 |
| CDM | 21206 | 95.62 | 3.61 | 0.36 | 0.41 |
| CDM | 7572 | 88.55 | 9.17 | 1.12 | 1.16 |
| CDM | 3417 | 83.79 | 12.38 | 1.82 | 2.01 |
| WDM | 28328 | 89.48 | 8.43 | 0.76 | 1.33 |
| WDM | 3793 | 79.25 | 14.79 | 2.14 | 3.82 |
| WDM | 5122 | 78.21 | 16.34 | 1.44 | 4.01 |
| Halos | ||||||||
| DM | 12340 | 11.57 | 8.02 | 0.92 | 48.72 | 37.41 | 10.46 | 3.41 |
| DM | 1950 | 5.66 | 2.82 | 0.16 | 72.10 | 16.41 | 7.23 | 4.26 |
| DM | 1550 | 5.10 | 2.89 | 0.15 | 64.45 | 27.36 | 8.19 | 0 |
| CDM | 21206 | 13.21 | 5.58 | 0.74 | 48.60 | 37.27 | 7.75 | 6.38 |
| CDM | 7572 | 24.57 | 8.24 | 2.02 | 29.32 | 30.78 | 14.05 | 25.85 |
| CDM | 3417 | 7.99 | 10.87 | 0.87 | 35.91 | 24.00 | 9.51 | 30.58 |
| WDM | 28328 | 24.56 | 13.75 | 3.37 | 38.04 | 36.55 | 12.05 | 13.36 |
| WDM | 3793 | 10.18 | 6.14 | 0.63 | 63.88 | 20.43 | 6.52 | 9.17 |
| WDM | 5122 | 16.08 | 9.31 | 1.50 | 49.57 | 38.54 | 11.21 | 0.68 |
5.3 Stellar Morphology
In this section, we will compare the stellar distribution of each galaxy across different DM simulations. Figure 8 presents the stellar morphology at three projections for the isolated G1 at . The main panels in this plot display the stellar distribution for the entire galaxy at a scale of approximately 55 kpc/h, while the additional sub-panels provide a closer view of the central regions of the galaxy (5 kpc/h). This allows for a detailed comparison between different stellar cores.
Figure 8 distinctly illustrates variations among three DM models, revealing that stars exhibit a more asymmetric structure in DM compared to a slightly symmetric shape seen in WDM and progressively a more symmetric shape in CDM, aligning well with theoretical predictions (Gao & Theuns, 2007; Schive et al., 2020). Notably, there is a significant discernible difference between the stellar distributions in WDM and DM. Despite appearing similar in scale in Figure 1, stars in WDM are actually more flattened than in DM. Upon closer examination in the subpanels with zoomed data, stars exhibit a spherical form in the DM core, but a more elliptical shape towards the center for both CDM and WDM. The plot effectively highlights the disparities between the core and cuspy centers, indicating that stars tend to concentrate in smaller cores in DM model, contrasting the cusps observed in CDM/WDM. Additionally, aside from having fewer stars than CDM and WDM, DM has a smaller proportion of stars in the central regions (see Table 4). This suggests a more extended stellar profile around the soliton, in line with recent discoveries of halos around some studied dwarfs (Chiti et al., 2021; Collins et al., 2021). These extended halos are inherent to DM, composed of excited states above the ground state soliton, as evidenced by the NFW form, as predicted by the DM simulations (Schive et al., 2014a). This reflects the fundamentally non-relativistic nature of DM. Moving to Figure 11, we observe WDM/CDM data for z 2.23. The results are comparable for both models when compared to the data from redshift z 5.56, showing an expected symmetry.
The additional sub-panels in Figure 8 prominently display the predicted asymmetry for DM, a distinctive feature not observed in the symmetric distributions of stars in WDM and CDM models (Schive et al., 2020). This asymmetry in DM arises from the solitonic characteristic ‘jumping’ behaviour (soliton oscillations) on the scale of the de Broglie wavelength , or the core radius (), owing to its standing wave nature (Li et al., 2021; Chiang et al., 2021). The jumps occur over a timescale of years for a dSph with 0.3 kpc and 10 km/s, values closely matching those for this galaxy (as referred in Tables 1 and 2 and check out Pozo et al. (2023) for a deeper explanation of ). Notably, this timescale is equivalent to the stellar crossing time of the soliton, . In essence, this means that stars ‘sense’ the movement of the soliton and respond accordingly, resulting in the observed asymmetry (Li et al., 2021; Chiang et al., 2021).
5.3.1 Stellar age vs metallicity
Metallicity plays a crucial role in unraveling the galaxy history. Here, we explore the relationship between the stellar metallicity and age. The corresponding data is displayed in the right columns of Figures 9 and 12. These figures illustrate the [Fe/H] versus stellar age profiles for G1 at the and for galaxies 1 and 3 at the . We chose to focus on these specific galaxies to emphasize the most pertinent data associated with instances where the potential presence of a core-halo structure has been identified.
We observe that in DM and WDM, the stellar metallicity exhibits a smoother decline as the stellar age increases. On the other hand, CDM displays a more pronounced breaks and a more discrete age vs. [Fe/H] profile, suggesting that its associated stars may have diverse origins. This disparity could arise from a combination of a few different factors. In CDM, the stellar cycle is more extended, with the first stars forming around z30 and continuing up to 9.8 billion years (as seen in Figure 9). In contrast, in WDM/DM, star formation begins around z13 (Mocz et al., 2019, 2020). Additionally, DM structures form earlier in CDM compared to WDM/DM, a consequence of the small-scale cut-off in the power spectrum.
Continuing with this idea, all three galaxies exhibit a peak of star formation around 9.6 billion years and 9.4 billion years, with the former being the predominant one for WDM/DM and the latter for CDM. Additionally, CDM is the only model with an additional small peak at 9.5 billion years. It is evident that all three models display a moderate to strong negative correlation between stellar age and metallicity. Similar trends and predictions are confirmed in Figure 12 at .
The observations in Figures 9 and 12 suggest that metal-rich stars tend to be more concentrated in the central regions of the galaxies, while metal-poor stars are located farther out, which is consistent with previous studies (Dietz et al., 2020; Santucci & otehrs, 2020). Additionally, the majority of stars are concentrated within distances compatible with the core-halo structure (r < 2 kpc, marked with a vertical dashed line). Of particular interest is that the galaxies where a compatible core-halo structure has been identified (DM at and WDM at ) exhibit a radially-offset region of stellar-density enhancement, which is especially pronounced in WDM at (as seen in Figure 12). Notably, WDM is the only case where this concentration of stars extends beyond the theoretical limit of the core-halo structure (r > 2 kpc). This peculiarity may be attributed to the filamentary environment observed for this galaxy, an environment that cannot be reproduced in the CDM model (as shown in Figure 1). In a future work we will study this in more detail. In summary, galaxies with a higher concentration of stars extending along their halos, rather than being centrally concentrated, seem to be more likely to develop the core-halo structure.
5.3.2 Spatial distribution of metallicity and age
Figures 10 and 13 investigate the correlation between the mean stellar metallicity and the stellar radii for galaxies from different DM models at both redshifts. The mean stellar age is depicted in the color bars.
The analysis of the correlation between stellar age and the spatial distribution of metallicity (as shown in Figures 10 and 13) reveals distinct differences among the three DM models. One notable distinction is that in CDM and WDM stars are relatively more metal-rich ([Fe/H] -1.0) compared to DM ([Fe/H] -1.25). Furthermore, CDM and WDM exhibit a smoother, continuous decline in metallicity from the center to the outer halo, whereas DM displays a more discrete profile, with noticeable drops in metallicity around 0.3 kpc and 0.5 kpc (as observed in the right panel of the first row in Figure 10). Regarding the spatial distribution of stellar age, WDM exhibits older stars concentrated in the center, while the mean age decreases towards the outer parts of the halo. This trend is opposite to what is observed for CDM and DM. However, it is important to note that for DM, the mean age is lower compared to the mean age observed for CDM, which aligns with the predicted stellar history between both models (Mocz et al., 2019, 2020). Tables 3 and 4 provide additional data regarding the percentage of metallicity and age in the stellar profiles of all the galaxies. These findings further contribute to our understanding of how different DM models influence the stellar properties of galaxies.
6 Discussion and Conclusions
The comparison between the core-halo structures of WDM and DM simulations with the CDM predictions reveals intriguing distinctions. In the case of WDM and DM, the core-halo structure appears to naturally stem from their respective power spectrum small-scale cut-off, whereas for CDM this structure is not anticipated. This notion could be further examined by tracking the stars of each galaxy across different epochs and observing their initial positions to scrutinize in more detail whether the stars appearing in the outer extended halos were born within their halo or in the outer filamentary structures. This observation, combined with other disparities between CDM predictions and observations, has spurred the investigation of alternative DM candidates like WDM and DM. Moreover, the fact that the stellar profiles in each DM scenario adhere to the predicted behaviours, with a cuspy profile for WDM and CDM, and a flat core structure for DM, supports the notion that stars trace DM. Figure 4 vividly illustrates that the DM, stellar, and the gaseous density profiles (especially stellar and DM) differ significantly in each DM scenario, and that the resulting stellar profile varies depending on the DM profile (cuspy/core discrepancy and core-halo structure). This underscores the potential significance of studying the properties of stars to achieve a deeper comprehension of the characteristics of DM.
Nonetheless, it is also true that in CDM and WDM both of the stellar and DM profiles appear to shift from a cuspy profile to cored one when decreasing the redshift (though never becoming as flat as DM predictions, which exhibit a better consistency with the observations (Sanchez Almeida et al., 2023)). This transformation from cuspy to cored profiles for CDM/WDM seems to align with the asserted impact of baryonic feedback in generating cores from originally cuspy halos. Feedback induces bursts of star formation and outflows, modifying the DM distribution and mitigating small-scale issues in CDM (Read & Gilmore, 2005; Amorisco et al., 2013; Chan et al., 2015), potentially resolving the cuspy-core problem. Furthermore, the simulations conducted by Mocz et al. (2019, 2020) were highly informative, demonstrating that baryonic feedback may not possess sufficient strength to convert cusps into cores for z>6, but it does open up possibilities for lower redshifts. This coincides with our findings, where the cusps observed for CDM/WDM at z5.56 seem to be alleviated around z2.23.
Certainly, the detection of the core-halo structure in WDM simulations, despite the absence of a soliton, is a topic of discussion. It seems that the smoothness of the cut-off in the initial dark matter power spectrum, which is likely more pronounced for WDM, may play a role in the formation of the core-halo structure in the stellar profiles. This suggests that the soliton itself may not be the sole factor determining the presence of the core-halo structure. However, it is worth noting that the presence of a soliton can be important as the transition feature in the DM and star profiles is a result of the power spectrum cut-off. The presence of a soliton may have a more notable impact on the early formation of the core-halo structure in DM (at ) compared to WDM (at ).
Indeed, the small-scale cut-off of the initial power spectrum appears to be the pivotal factor in determining the core-halo structure, as it leaves its mark on both the DM and the stellar profiles. The resemblance of the large-scale structures between the WDM and DM, despite their distinct physics on smaller scales, lends support to this notion. It is worth noting that the underlying cause of the cut-off differs for WDM and DM, with WDM exhibiting a thermal free-streaming cut-off, while DM featuring a quantum mechanical cut-off. However, this distinction does not seem to significantly impact larger scales where the cut-off length scale is imprinted on the profiles. The simulation data presented by Mocz et al. (2020) demonstrates that both of the WDM and DM yield core-halo stellar profiles, which appear to be at odds with CDM predictions. Further analysis is warranted to validate this finding and to probe the hypothesis that the core-halo structure arises from the small-scale cut-off of the initial power spectrum.
On the other hand, the discrepancies found in the stellar morphology, age, and metallicity suggest the notion of a less concentrated and non-cuspy inner halo as a prerequisite for the observed core-halo structure in local dwarfs. This idea gains further support when we observe the higher concentration of stars in CDM/WDM models compared to DM, which exhibit a lower proportion of stars in their halo centers, indicating more extended stellar profiles. Furthermore, the spatial distribution of metallicity and age indicates that the core-halo structure emerges in galaxies where the greater concentration of stars is not extremely close to the center, a condition that appears only feasible in a filamentary environment, which is challenging to replicate for CDM. Finally, the solitonic stellar imprint is substantiated by the observed asymmetry in the core of G1 in the DM scenario, as a consequence of the previously analyzed solitonic random walk (Li et al., 2021; Chiang et al., 2021), where stars are thought to “sense” the movement of the soliton and respond accordingly, resulting in the observed asymmetry, in stark contrast to CDM/WDM outcomes. This further strengthens the idea of distinct stellar behaviors for each DM model. Quantifying this effect with the deepest JWST images of different dwarfs at various redshifts could yield valuable insights. Additionally, a more detailed analysis of this soliton ‘jumping’ phenomenon could be carried out with data from lower redshifts, an area we plan to explore in future studies by tracking the motion of stars around the soliton across different redshifts.
We conclude that CDM is under significant tension with core-halo observations. Moreover, there are several compelling reasons to favor DM over WDM/CDM. Firstly, as highlighted previously, DM seems to align better with observations in terms of the presence and structure of core-halo profiles in galaxy halos, as opposed to CDM. Additionally, DM predicts a granularity in the lensed halos Amruth et al. (2023) that has not been observed in WDM simulations. Furthermore, the lack of detection of WDM particles in laboratory searches is a substantial drawback for this model. In contrast, DM posits light bosons that would necessitate a much larger laboratory scale of 1 kpc for detection, making it challenging to rule out this possibility based solely on laboratory experiments. Overall, the combination of better agreement with observations and the lack of definitive evidence against it makes DM a promising candidate for dark matter, and a more viable alternative to WDM/CDM.
Acknowledgements
We warmly acknowledge Douglas Finkbeiner for very fruitful conversations. A.P. is grateful for the continued support of the DIPC graduate student program as well as the Center for Astrophysics | Harvard & Smithsonian for their warm hospitality at the beginning of this project. R.E. acknowledges the support from grant numbers 21-atp21-0077, NSF AST-1816420, and HST-GO-16173.001-A as well as the Institute for Theory and Computation at the Center for Astrophysics. We are grateful to the supercomputer facility at Harvard University where most of the simulation work was done. G.S. is grateful to the IAS at HKUST for their generous support. This work has been supported by the Spanish project PID2020-114035GB-100 (MINECO/AEI/FEDER, UE). P.M. acknowledges this work was in part performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, Lawrence Livermore National Security, LLC.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
Appendix A Stellar Morphology at
Here we present the equivalent plots for data at from the main text, for . This supplementary analysis further enriches our understanding of the core-halo structure in the system under investigation. It is valuable to explore how these parameters interact and correlate at different redshifts, providing a more comprehensive view of the underlying dynamics.

### v2

@@
-Galaxy formation with Wave/Fuzzy Dark Matter: The core-halo structure
-Abstract
+Galaxy formation with wave/fuzzy dark matter: The core-halo structure and the solitonic imprint
+11email: alvaro.pozolarrocha@bizkaia.eu; tom.j.broadhurst@gmail.com; 22institutetext: DIPC, Basque Country UPV/EHU, E-48080 San Sebastian, Spain 33institutetext: Center for Astrophysics Harvard Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA 44institutetext: Department of Astrophysical Sciences, Princeton University, 4 Ivy Lane, Princeton, NJ, 08544, USA 55institutetext: Lawrence Livermore National Laboratory, 7000 East Ave, Livermore, CA 94550, USA 66institutetext: Ikerbasque, Basque Foundation for Science, Bilbao, E-48011, Spain 77institutetext: Hong Kong University of Science and Technology, Institute for Advanced Study and Department of Physics, IAS TT WF Chao Foundation Professor, Hong Kong 88institutetext: Energetic Cosmos Laboratory, Nazarbayev University, Nursultan, Kazakhstan 99institutetext: Paris Centre for Cosmological Physics, APC, AstroParticule et Cosmologie, Université de Paris, CNRS/IN2P3, CEA/lrfu,Université Sorbonne Paris Cité, 10, rue Alice Domon et Leonie Duquet, 75205 Paris CEDEX 13, France Emeritus 1010institutetext: Physics Department LBNL, University of California at Berkeley CA 94720 Emeritus 1111institutetext: Department of Physics, Kavli Institute for Astrophysics and Space Research, Massachusetts Institute of Technology, Cambridge, MA 02139, USA 1212institutetext: Black Hole Initiative at Harvard University, 20 Garden Street, Cambridge, MA 02138, USA
+Galaxy formation with wave/fuzzy dark matter: The core-halo structure and the solitonic imprint
 Dark matter-dominated cores have long been claimed for the well-studied local group dwarf galaxies.
 More recently, extended stellar halos have been uncovered around several of these dwarfs through deeper imaging and spectroscopy.
-Such core-halo structures are not a feature of conventional cold dark matter (CDM), based on collisionless particles where smooth, scale-free profiles are predicted.
-In contrast, smooth and prominent dark matter cores are predicted for Warm and Fuzzy/Wave Dark Matter (WDM/DM) respectively.
-The question arises to what extent the visible stellar profiles should reflect this dark matter core structure.
-Here we compare cosmological hydrodynamical simulations of CDM, WDM DM, aiming to predict the stellar profiles for these three DM scenarios.
-We show that cores surrounded by extended halos are distinguishable for WDM and DM, with the most prominent cores in the case of DM, where the stellar density is enhanced in the core due to the presence of the relatively dense soliton.
-Our analysis demonstrates that such behavior does not appear in CDM, implying that the small-scale cut-off in the power spectrum present for WDM and DM provides a core-halo transition.
-Consequently, we estimate the mass of the DM particle at this core-halo transition point.
-Furthermore, we observe the anticipated asymmetry for DM due to the soliton’s random walk, a distinctive characteristic not found in the symmetric distributions of stars in Warm and CDM models.
-keywords:
-Dark Matter, Galaxies: dwarfs, Galaxies: kinematics and dynamics1 Introduction
-Dark Matter is understood to be non-relativistic, i.e.
-cold, as required to form galaxies gravitationally and for explaining the spectrum of CMB fluctuations in detail (Collaboration et al., 2018).
-However, the standard heavy particle interpretation for Cold Dark Matter (CDM) faces a stringent laboratory absence of any new particle signature (Aprile et al., 2018; Collaboration, 2020) and several inconsistencies have emerged between CDM predictions and the properties of dwarf galaxies (Moore, 1994; de Blok, 2010; Marsh & Pop, 2015; Klypin et al., 1999; Safarzadeh & Loeb, 2021).
-Warm Dark Matter (WDM) is defined to be relativistic initially with a keV-scale mass that is sufficiently redshifted to become non-relativistic before galaxy formation, generating smooth cores and suppressing formation of low-mass galaxies.
-Alternatively, for very light bosons of , an inherently non-relativistic possibility of dark matter as a Bose-Einstein condensation (Widrow & Kaiser, 1993; Hu et al., 2000), originally termed Fuzzy Dark Matter (FDM), has been emerged by the first simulations in this context, which predict prominent solitonic standing wave cores surrounded by pervasive interference on the de Broglie wavelength scale within all galaxies and filaments (Schive et al., 2014a; Mocz et al., 2017; Veltmaat et al., 2018; Hui, 2020).
-This was then termed as ‘Wave Dark Matter’, DM, (Schive et al., 2014a; Hui et al., 2021).
-While DM predicts almost identical large-scale structure to CDM in terms of the network of filaments and clusters (Schive et al., 2014a; Mocz et al., 2017), the small-scale structure of DM halos is very different in DM paradigm, as DM cannot be confined to less than the de Broglie scale.
-Consequently, in DM, with a light boson mass of , the dwarf galaxy formation is suppressed and there is also a prominent solitonic core (Schive et al., 2014a, b; Mocz et al., 2017; Veltmaat et al., 2018; Niemeyer, 2020) in the galaxy.
+Such core-halo structures (inner flat core and a characteristic asymptotic outer halo profile) are not a feature of conventional cold dark matter (CDM).
+In contrast, smooth and prominent dark matter cores are predicted for wave/fuzzy dark matter (DM).
+The question arises as to what extent the visible stellar profiles should reflect this dark matter core structure.
+Here we compare cosmological hydrodynamical simulations of CDM, “WDM” (model used as a proxy for DM) DM, aiming to predict the stellar profiles for these three DM scenarios.
+We show that cores surrounded by extended halos are distinguishable for DM, where the stellar density is enhanced in the core due to the presence of the relatively dense soliton.
+Our analysis demonstrates that, in our simulations, a distinctive core-halo structure does not appear in the case of CDM in the DM, gas, or stars.
+Whereas we do find a core-halo transition for DM, gas, and stars for DM, and the scale of this transition is in line with the predicted core radius set by the soliton scale anticipated for the adopted boson mass of 2.5eV.
+The presence of a core-halo structure in the stellar profile for Galaxy 1 for DM is visible for the most massive and the first galaxy to form in the simulation.
+Clearly, further simulations are needed to establish how strict this possible relationship is between the DM and stellar core-halo profile as a potential observational discriminator.
+Furthermore, we observe the anticipated asymmetry for DM due to the soliton’s motion (jumping and random walk), a distinctive characteristic not found in the symmetric distributions of stars in the warm and CDM models.
+Key Words.:
+cosmology – dark matter – galaxies1 Introduction
+Dark matter is understood to be nonrelativistic, i.e., cold, as required to form galaxies gravitationally and to explain the spectrum of CMB fluctuations in detail (Aghanim et al.
+2018).
+However, the standard heavy particle interpretation for cold dark matter (CDM) faces a stringent laboratory absence of any new particle signature (Aprile et al.
+2018; Sirunyan et al.
+2020) and several inconsistencies have emerged between CDM predictions and the properties of dwarf galaxies (Moore 1994; de Blok 2010; Marsh & Pop 2015; Klypin et al.
+1999; Safarzadeh & Loeb 2021).
+Warm dark matter (“WDM”) is defined to be relativistic initially with a kiloelectron volt-scale mass, the 1.4keV mass scale used here for the “WDM” simulation induces the power spectrum suppression, due to early relativistic free streaming, but becomes nonrelativistic later via the cosmological expansion, so that the “WDM” simulation is closely approximated by N-body simulations.
+The “WDM” simulations analyzed here were performed using a DM initial power spectrum, without including the dynamical quantum potential.
+This approximation reproduces the small-scale cutoff expected for a thermal relic “WDM” particle of mass 1.4 keV, corresponding to the boson mass of eV used in the DM simulations of Mocz et al.
+(2020).
+Alternatively, for very light bosons of , an inherently nonrelativistic possibility of dark matter as a Bose-Einstein condensation (Widrow & Kaiser 1993; Hu et al.
+2000), originally termed fuzzy dark matter (FDM), emerged from the first simulations in this context, which predict prominent solitonic standing wave cores surrounded by pervasive interference on the de Broglie wavelength scale within all galaxies and filaments (Schive et al.
+2014a; Mocz et al.
+2017; Veltmaat et al.
+2018; Hui 2020).
+This was then called wave/fuzzy dark matter (DM) (Schive et al.
+2014a; Hui et al.
+2021).
+While DM predicts large-scale structure that are almost identical to CDM in terms of the network of filaments and clusters (Schive et al.
+2014a; Mocz et al.
+2017), the small-scale structure of DM halos is very different in the DM paradigm, as DM cannot be confined to less than the de Broglie scale.
+Consequently, in DM, with a light boson mass of , the dwarf galaxy formation is suppressed and there is also a prominent solitonic core (Schive et al.
+2014a, b; Mocz et al.
+2017; Veltmaat et al.
+2018; Niemeyer 2020) in the galaxy.
 On these scales, self-gravity balances the effective pressure from the uncertainty principle in the ground state at the de Broglie wavelength.
 Crucially, smaller galaxies are predicted to have wider cores of lower density because the soliton is larger at lower momentum, which we test here.
-Many independent claims of dark matter and stellar cores in deep images and spectroscopy are now available for the well-studied, Local Group dwarf spheroidal galaxies, dSph, providing one of the main sources of tension for CDM for which the lack of any predicted cores reflects the scale-free nature of simplest, collisionless DM scenario.
-More recently the presence of extended stellar halos has been uncovered in several local dwarfs, independently, with deep imaging and spectroscopy (Chiti et al., 2021; Collins et al., 2021).
-A joint analysis of the available data now suggests that such a core-halo structure with a sharp density transition between the core and the halo is a common feature of both dSph galaxies and also the smaller new class of UFG (Ultra faint galaxies) (Pozo et al., 2020, 2023).
-This core-halo behaviour motivates the examination of hydrodynamical simulations with a comparison of the differing DM contenders as a possible distinguishing feature.
+Many independent claims of dark matter and stellar cores in deep images and spectroscopy are now available for the well-studied Local Group dwarf spheroidal galaxies, dSph, providing one of the main sources of tension for CDM for which the lack of any predicted cores reflects the scale-free nature of the simplest collisionless DM scenario.
+More recently the presence of extended stellar halos has been uncovered in several local dwarfs, independently, via deep imaging and spectroscopy (Chiti et al.
+2021; Collins et al.
+2021).
+A joint analysis of the available data now suggests that such a core-halo structure with a sharp density transition between the core and the halo is a common feature of dSph galaxies and also of the smaller new class of ultra-faint galaxies (UFGs) (Pozo et al.
+2024, 2023).
+This core-halo behavior (Chan et al.
+2020a) motivates the examination of hydrodynamical simulations with a comparison of the differing DM contenders as a possible distinguishing feature.
 For cold dark matter (CDM), the scale-free formation of DM structures is predicted to extend to arbitrarily small scales, reaching lower halo masses with higher DM concentrations.
-In contrast, small-scale suppression of the power spectrum is inherent to both DM and WDM.
-In the case of DM, this suppression limits the minimum scale of structure to below the de Broglie length, which is determined by the boson mass.
+In contrast, small-scale suppression of the power spectrum is inherent to both DM and “WDM”.
+In the case of DM, this suppression limits the minimum scale of structure to above the de Broglie length, which is determined by the boson mass.
 Consequently, DM structures below are suppressed for a choice of boson masses of eV, motivated by the observed kpc scale of dwarf galaxy cores.
-This limitation has a direct consequence for the formation timescales between CDM and DM, resulting in a delay in galaxy formation for DM compared to CDM.
-In both cases, we observe that the formation of the first "filamentary" structures occurs earlier in CDM, where they are comprised of low-mass subhaloes.
-In contrast, in DM/WDM, there is no fragmentation along the filaments due to the small-scale cutoff of the power spectrum.
-Instead, dark matter is distributed more continuously along these filaments.
-The main difference between warm dark matter (WDM) and DM is that WDM exhibits a clear caustic structure in the distribution of dark matter, while DM is characterized by wave interference patterns that manifest as fringes in simulations.
-These fringes are now understood to generate turbulent density structures within halos, resulting from the superposition of numerous plane waves encoding the velocity dispersion in the halo (Mocz et al., 2020).
-In DM, these filaments persist for longer periods, as the gravitational attraction from the early formation of the first halos in cold dark matter (CDM) tends to disrupt them earlier.
-Moreover, the prolonged lifetimes of these filaments in wave dark matter favor the formation of a greater number of stars within them, leading to significant differences in the location and extent of stellar profiles in CDM and DM galaxies.
-This also causes baryonic objects to appear more diffuse or smoothed compared to CDM (Mocz et al., 2019, 2020), indicating that they could serve as excellent tracers of dark matter in DM.
+This limitation has a direct consequence for the formation timescales between CDM and DM, and results in a delay in galaxy formation for DM compared to CDM.
+In both cases, we observe that the formation of the first filamentary structures occur earlier in CDM, where they are comprised of low-mass subhalos.
+In contrast, in DM/“WDM” there is no fragmentation along the filaments, due to the small-scale cutoff of the power spectrum.
+Instead, dark matter is distributed more continuously along these filaments (May & Springel 2023; Dome et al.
+2023).
+The main difference between “WDM” and DM is that “WDM” exhibits a clear caustic structure in the distribution of dark matter, while DM is characterized by wave interference patterns that manifest as fringes in simulations.
+These fringes are now understood to generate turbulent density structures within halos, resulting from the superposition of numerous plane waves encoding the velocity dispersion in the halo (Mocz et al.
+2020).
+In DM these filaments persist for longer periods as the gravitational attraction from the early formation of the first halos in cold dark matter (CDM) tends to disrupt them earlier.
+Moreover, the prolonged lifetimes of these filaments in wave/fuzzy dark matter favor the formation of a greater number of stars within them, leading to significant differences in the location and extent of stellar profiles in CDM and DM galaxies.
+This also causes baryonic objects to appear more diffuse or smoothed compared to CDM (Mocz et al.
+2019, 2020), indicating that they could serve as excellent tracers (as visible objects that could be used to observe the underlying DM structures) of dark matter in DM.
 On the scale of solitons, the quantum pressure in DM can become sufficiently strong to counteract the self-gravity of the dark matter.
-This results in the formation of prominent solitonic standing waves at the base of each virialized potential, characterized by a spherical soliton core of a few kiloparsecs at the center of every dark matter halo (Schive et al., 2014a), in contrast to the much denser cusps observed in CDM (Navarro & Frenk, 1996) and WDM (Mocz et al., 2019).
-In dSph galaxies it is known that stellar motions are dominated by the dark matter gravitational potential and so we expect some visible differences for stars in these galaxies depending on the nature of dark matter.
-In addition, the formation and early evolution of galaxies and filaments will also be sensitive to the nature of DM, arising in particular from the suppressed form of the high-k power spectrum for the WDM/DM models.
-The wave behaviour and the quantum pressure of DM provide different virialized DM halo structures and prominent soliton cores, unlike the smooth cores of WDM.
-While for CDM and WDM cuspy profiles are expected (Navarro & Frenk, 1996; Lovell et al., 2014), DM simulations of merging DM haloes clearly show the formation of soliton structures at the de Broglie wavelength scale (Schive et al., 2014a; Mocz et al., 2017; Schwabe et al., 2016), and are prominent in terms of the core density, standing above the surrounding DM halo.
-This is in contrast to the early fragmentation of filaments and cuspy haloes in CDM (Mocz et al., 2017, 2019, 2020).
+This results in the formation of prominent solitonic standing waves at the base of each virialized potential, characterized by a spherical soliton core of a few kiloparsecs at the center of every dark matter halo (Schive et al.
+2014a), in contrast to the much denser cusps observed in CDM (Navarro & Frenk 1996).
+It is remarkable how well the soliton profile of DM aligns with the empirically well-established Burkert profile, as demonstrated in the extensively studied Fornax dwarf spheroidal galaxy Schive et al.
+(2014a).
+However, the scaling of the core radius with galaxy mass reported by Burkert (2020) from profile fitting to low- and intermediate-mass galaxies exhibits a slope that is opposite in sign to that predicted by DM, as highlighted by Burkert (2020).
+Moreover, the relatively large cores inferred for intermediate-mass galaxies are based on HI gas dynamics.
+Noncircular gas motions have been recognized as a significant contribution to underestimated rotation curves (Oman et al.
+2019; Downing & Oman 2023), together with resolution limitation for HI (Swaters et al.
+2000) and the finite thickness of HI disks that all act to bias downward HI-based rotation speed measurements, as demonstrated using CDM-based hydro simulations with spurious large cores (Roper et al.
+2023).
+It is important to examine the claimed scaling of galaxy mass and core radius with independent stellar-based dynamics, though this remains challenging due to the typically low stellar surface brightnesses.
+In dSph galaxies it is known that stellar motions are dominated by the dark matter gravitational potential, and so we expect some visible differences for stars in these galaxies depending on the nature of dark matter.
+In addition, the formation and early evolution of galaxies and filaments is also sensitive to the nature of DM, arising in particular from the suppressed form of the high-k power spectrum for the “WDM”–DM models.
+The wave behavior and the quantum pressure of DM provide different virialized DM halo structures and prominent solitonic cores, unlike the CDM–WDM cuspy profiles (without inner cores) expectations(Navarro & Frenk 1996) (even smooth cores of “ WDM are claimed (Lovell et al.
+2014; Macciò et al.
+2012; Bode et al.
+2001)), DM simulations of merging DM halos clearly show the formation of soliton structures at the de Broglie wavelength scale (Schive et al.
+2014a; Mocz et al.
+2017; Schwabe et al.
+2016), and are prominent in terms of the core density, lying above the surrounding DM halo.
+This is in contrast to the early fragmentation of filaments and cuspy halos in CDM (Mocz et al.
+2017, 2019, 2020).
 Mocz et al.
-(2019, 2020) demonstrated that baryonic feedback does not have a significant effect in halos of to for redshifts , in the sense of being unable to soften the cuspy profiles of CDM/WDM to produce cores.
-Consequently, neither CDM/WDM are able to explain the origin of the claimed cores from the dynamical studies of dSph galaxies (Read & Gilmore, 2005; Amorisco et al., 2013).
-In this paper we compare the galaxy formation process between CDM, WDM, and DM, in the high-resolution cosmological simulations by Mocz et al.
+(2019, 2020) postulated that baryonic feedback may not have a significant effect in halos of to for redshifts , in the sense of being unable to soften the cuspy profiles of CDM to produce cores.
+Consequently, CDM should not be able to explain the origin of the claimed cores from the dynamical studies of dSph galaxies (Read & Gilmore 2005; Amorisco et al.
+2013).
+In this paper we compare the galaxy formation process between CDM, “WDM”, and DM, in the high-resolution cosmological simulations by Mocz et al.
 (2019, 2020, 2017).
-Here, we will consider the distribution of stars modelled in these simulations, where in the WDM and DM contexts these simulations have already reported stars to be born along dense dark matter filaments, tracing dark matter, and highlighted as “a smoking gun” signature of DM (Mocz et al., 2019).
-Here we will further examine these lowest redshifts, focusing on the question of core-halo structure for comparison with new deep imaging and spectroscopic data for the Milky Way and other local group dwarfs.
-For these objects, the data now indicates non-smooth profiles in these dark matter-dominated dwarfs as this stellar structure may potentially have important implications for the nature of dark matter.
-Moreover, we will look into the different evolution of the stellar populations and the impact of the three DM models on their spatial distribution.
-The structure of this paper is as follows: firstly, in sections 2 and 3, we describe the radial structure and internal dynamics of the DM halo for comparison with the data.
-Next, in section 4, we present the simulation setup.
-In section 5, we focus on presenting all the results of the work, with particular emphasis on the main finding related to the detection of the core-halo structure.
-Finally, in section 6, we discuss our conclusions regarding the non-detection of the core-halo structure for CDM and its implications.
-2 The Wave Dark Matter Halo
-Ultralight bosons, such as Axions, have been considered as an ingredient in describing the DM (Widrow & Kaiser, 1993; Hu et al., 2000).
+Here we consider the distribution of stars modeled in these simulations, where in the “WDM” and DM contexts these simulations have already reported stars to be born along dense dark matter filaments, tracing dark matter, and highlighted as “a smoking gun” signature of DM/“WDM” (Mocz et al.
+2019).
+Here we further examine these lowest redshifts, focusing on the question of core-halo structure for comparison with new deep imaging and spectroscopic data for the Milky Way and other local group dwarfs.
+The structure of this paper is as follows.
+In sections 2 and 3 we describe the radial structure and internal dynamics of the DM halo for comparison with the data.
+Next, in section 4 we present the simulation setup.
+In section 5 we focus on presenting all the results of the work, with particular emphasis on the main finding related to the detection of the core-halo structure.
+Finally, in section 6 we discuss our conclusions regarding the nondetection of the core-halo structure for CDM and its implications.
+The boson masses used in this work are in strong tension with current constraints in the literature (for a discussion of these constraints, see section 6).
+2 The wave/fuzzy dark matter halo
+Ultralight bosons, such as ALPs (axion-like particles), have been considered as an ingredient in describing the DM (Widrow & Kaiser 1993; Hu et al.
+2000).
 In the absence of any self-interactions, the boson mass is the only free parameter that describes the DM.
-For a sufficiently light boson mass, the de-Broglie wavelength exceeds the mean free path leading the Axion to satisfy the ground state condition for a Bose-Einstein condensation.
-In the comoving coordinate, the governing equation for Axion reads as:
-where is the wave function, is the gravitation potential and (1/(1+z)) is the cosmological scale factor.
-The system is normalized to the time scale , and to the scale length , with , where is the current density parameter (Widrow & Kaiser, 1993).
-Recently, it has proved possible with advanced GPU computing to make reliable, high dynamic range cosmological simulations that solve the above equations, (Schive et al., 2014a; Schwabe et al., 2016; Mocz et al., 2017; May & Springel, 2021).
-These evolve to produce large-scale structures indistinguishable from CDM, but with virialized halos characterized by a solitonic core in the ground state that naturally explains the dark matter-dominated cores of dSph galaxies (Schive et al., 2014b).
-Another important feature arising from the simulations is that the central soliton is surrounded by an extended halo with a “granular” texture on the de-Broglie scale, due to interference of excited states, but which when azimuthally averaged follows the Navarro-Frenk-White (NFW) density profile (Navarro & Frenk, 1996; Woo & Chiueh, 2009; Schive et al., 2014a, b).
-The fitting formula for the density profile of the solitonic core in a DM halo is obtained from cosmological simulations (Schive et al., 2014a, b) as:
+For a sufficiently light boson mass, the de-Broglie wavelength exceeds the mean free path, which leads the ALP to satisfy the ground state condition for a Bose-Einstein condensation.
+The governing equation for DM is the Schrödinger-Poisson (SP) equation, which in comoving coordinates reads as:
+where is the wave function (see Schive et al.
+(2014a) for the explanation of the interpretation of the wave function, originally introduced by Widrow & Kaiser (1993), to describe the mean field behavior of the wave function in this context), is the gravitation potential and =(1/(1+z)) is the cosmological scale factor, where z is the redshift.The comoving length is normalized to ( , the time normalized to and the wave function normalized to , where is the present Hubble parameter, is the matter density parameter and the background mass density, (Schive et al.
+2014b).
+Finally, the physical interpretation of the wave function in DM is that the density is given by .
+Recently, it has become possible to perform high dynamic range cosmological simulations that solve the above equations (Schive et al.
+2014a; Schwabe et al.
+2016; Mocz et al.
+2017; May & Springel 2021), where GPU computing has allowed an efficient adaptive mesh refinement (Schive et al.
+2014a).
+These evolve to produce large-scale structures indistinguishable from CDM, but with virialized halos characterized by a solitonic core in the ground state that naturally explains the dark matter-dominated cores of dSph galaxies, previously analyzed in many dSph and UFD of the Local Group (Pozo et al.
+2024, 2023, 2021, 2022; Schive et al.
+2014b) .
+Another important feature arising from the simulations is that the central soliton is surrounded by an extended halo with a granular texture on the de-Broglie scale, due to interference of excited states, but which when azimuthally averaged follows the Navarro-Frenk-White (NFW) density profile (Navarro & Frenk 1996; Woo & Chiueh 2009; Schive et al.
+2014a, b).
+The fitting formula for the density profile of the solitonic core in a DM halo is obtained from cosmological simulations (Schive et al. 2014a, b) as
 where is the boson mass and is the solitonic core radius.
-The latter scales with the product of the galaxy mass and boson mass, obeying the following scaling relation which has been derived from simulations (Schive et al., 2014b):
-Beyond the soliton, at radii larger than a transition radius (), the simulations also reveal that the halo roughly looks like NFW in shape, presumably reflecting the non-relativistic nature of condensates beyond the de Broglie scale, and therefore the total density profile can be written as:
-In detail, the scale radius of the solitonic solution, which represents the ground state of the Schrodinger-Poisson equation, is related to the size of the halo through the uncertainty principle.
-From cosmological simulations, the latter is found to hold non-locally, relating a local property with a global one (for more details we refer to Schive et al.
-(2014b)).
-3 Dynamical Model of Galaxies in Wave Dark Matter Halo
-Classical dSph galaxies are known to be dominated by DM with stars being treated as the tracer particles (Gregory et al., 2019; McConnachie & Irwin, 2006; McConnachie et al., 2006; Kang & Ricotti, 2019) moving along the gravitational potential generated by the DM halo.
+The latter scales with both the halo mass and the boson mass, obeying the following scaling relation which was derived from the simulations (Eq.
+7 (Schive et al.
+2014b))
+where is the halo mass and .
+is defined as the matter density parameter as a function of redshift, normalized to the critical density at that time.
+Beyond the soliton, at radii larger than a transition radius (), the simulations also reveal that the halo roughly resembles NFW in shape, presumably reflecting the nonrelativistic nature of condensates beyond the de Broglie scale, and therefore the total density profile can be written as
+where is chosen such that the inner solitonic profile matches the outer NFW-like profile at approximately , and is the scale radius.
+In detail, the scale radius of the solitonic solution, which represents the ground state of the Schrödinger-Poisson equation, is related to the size of the halo through the uncertainty principle.
+3 Dynamical model of galaxies in wave/fuzzy dark matter halo
+Classical dSph galaxies are known to be dominated by DM with stars being treated as the tracer particles (Gregory et al.
+2019; McConnachie & Irwin 2006; McConnachie et al.
+2006; Kang & Ricotti 2019) moving along the gravitational potential generated by the DM halo.
 In this context, the corresponding velocity dispersion profile can be estimated by solving the spherically symmetric Jeans equation:
-where is the DM halo mass obtained by integrating the spherically symmetric density profile in Eq.
-(5), is the anisotropy parameter (see (Binney & Tremaine, 2008), Equation (4.61)), and is the stellar density profile as given by:
+Here is the DM halo mass obtained by integrating the spherically symmetric density profile in Eq.
+(5), describes the radial velocity dispersion, is the anisotropy parameter (see (Binney & Tremaine 2008), Equation (4.61)), and is the stellar density profile as given by
 where
 Here, is the 3D scale radius of the stellar halo corresponding to the central stellar density, is the normalization of at the transition radius and the transition radius, , is the point where the solitonic structure ends and the halo begins at the juncture of the core and halo profiles.
+We employ Eq.
+(7) to identify the core-halo structure in the analyzed DM stellar profiles, given its high efficiency in fitting the observed core-halo transition in local group galaxies (Pozo et al.
+2024) and as it is the only profile capable of explaining this observed transition between the two regimes—inner core and outer halo—together.
+This approach is used exclusively for DM, as it is unsuitable for non-DM models, where the presence of a soliton is not expected (Pozo et al.
+2022, 2023).
+For CDM–“WDM”, there is no equation that describes both regimes simultaneously, as such a structure is not expected in their DM or stellar profiles.
+However, for the DM profile, we can adjust eq.
+(5) , which defines the core-halo structure (described in Section 5.3) within the DM model, to align with what we expect for stars, as indicated by eq.
+(7).
+For CDM/“WDM” we used the isothermal equation from Binney & Tremaine (2008) to describe the observed stellar cored profiles:
+Here is the mass density at a radial distance r from the center of the galaxy.
+represents the central density, or the density at r = 0.
+is the core radius, which defines the scale at which the density profile transitions from a flat core to a declining halo.
+The isothermal profile is commonly used to describe galaxies with a central region where the density remains approximately constant (the core) and a halo where the density decreases with distance, following approximately at large .
+However, this decrease is incompatible with the core-halo structure defined in Section 5.3 and in eq.
+(5), which predicts a transition point where the density should start declining according to an asymptotic profile of .
+Finally, the predicted 2D stellar density profile can be projected along the line of sight, and compared with the observations, as
 4 Simulation setup
@@
 The cosmological parameters correspond to the Planck results.
-The primary objective is to conduct a thorough comparative analysis of the stellar properties between the DM simulations and the ones stemming from both CDM and WDM simulations, spanning an extended duration down to .
-In the case of the WDM simulations, a particle resolution of was utilized, employing the identical hydrodynamical setup as the DM simulations.
-It is noteworthy that while the CDM simulation lacks any form of cutoff in its power spectrum, both the DM and WDM simulations adopt an initial power spectrum characterized by exponential suppression.
+The primary objective is to conduct a thorough comparative analysis of the stellar properties between the DM simulations and the ones stemming from both CDM and “WDM” simulations, spanning an extended duration down to .
+In the case of the “WDM” simulations, a particle resolution of was utilized, employing the identical hydrodynamical setup as the DM simulations.
+It is noteworthy that while the CDM simulation lacks any form of cutoff in its power spectrum, both the DM and “WDM” simulations adopt an initial power spectrum characterized by exponential suppression.
 Baryons (gas and star formation) are included in the simulation, and couple to the dark matter gravitationally.
-The baryonic material experiences primordial and metal-line cooling, chemical enrichment, stochastic star formation with a density threshold of , supernova feedback via kinetic winds, and instantaneous uniform reionization at (Vogelsberger et al., 2013; Torrey et al., 2014; Pillepich et al., 2017; Springel et al., 2018), as has been used in the Illustris and Illustris-TNG projects.
-The stellar feedback model has previously been tuned to match key observables with CDM simulations (Vogelsberger et al., 2013).
+The baryonic material experiences primordial and metal-line cooling, chemical enrichment, stochastic star formation with a density threshold of , supernova feedback via kinetic winds, and instantaneous uniform reionization at (Vogelsberger et al.
+2013; Torrey et al.
+2014; Pillepich et al.
+2018a; Springel et al.
+2018), as has been used in the Illustris and Illustris-TNG projects.
+The stellar feedback model has previously been tuned to match key observables with CDM simulations (Vogelsberger et al.
+2013).
 Whether FDM requires significant re-tuning of efficiency factors is left for future study.
-The subgrid model for feedback is meant to describe the effects of Type II supernovae (SNII), and uses the local star formation rate to set the mass loading of stellar winds driven by the energy available through SNII (Vogelsberger et al., 2013; Pillepich et al., 2018).
+The subgrid model for feedback is meant to describe the effects of Type II supernovae (SNII), and uses the local star formation rate to set the mass loading of stellar winds driven by the energy available through SNII (Vogelsberger et al.
+2013; Pillepich et al.
+2018b).
 The simulation also keeps track of metallicty, as described in Torrey et al.
@@
 5.1 First illustration of density profiles
-Before delving into the intricacies of comparing various simulation types, we provide an initial depiction of the projected density profiles for CDM, WDM, and DM at two different redshifts.
+Before delving into the intricacies of comparing various simulation types, we provide an initial depiction of the projected density profiles for CDM, “WDM”, and DM at two different redshifts.
 In Figure 1, we present the projected density profiles for both DM and stars.
-The first two rows illustrate CDM profiles, while the last two rows depict the corresponding profiles for WDM/DM.
+The first two rows illustrate CDM profiles, while the last two rows depict the corresponding profiles for “WDM”/DM.
 In each row, the left(right) panel presents ().
 Mocz et al.
-(2019) showed that WDM and DM exhibit remarkably similar evolutionary patterns across various redshifts, implying a resemblance in the development of large-scale structures.
+(2019) showed that “WDM” and DM exhibit remarkably similar evolutionary patterns across various redshifts, implying a resemblance in the development of large-scale structures.
 This stands in contrast to CDM.
-Consequently, in Figure 1 we exclusively focus on presenting the WDM results.
-Figure 1 suggests slight discrepancies between galaxies in these three distinct DM scenarios.
+Consequently, in Figure 1 we exclusively focus on presenting the “WDM” results.
+Figure 1 and 2 suggest slight discrepancies between galaxies in these three distinct DM scenarios.
 First, in the context of CDM, filaments rapidly collapse into compact halos, whereas in the case of DM, these filaments retain a smoother configuration without giving rise to small galaxies.
 Second, in the CDM scenario, stars seem to exhibit greater concentration with irregular structures.
-Conversely, in the case of DM and WDM, a flatter shape is evident, accompanied by galaxies possessing extended stellar halos.
+Conversely, in the case of DM and “WDM”, a flatter shape is evident, accompanied by galaxies possessing extended stellar halos.
 This imparts a visual impression that stars are positioned farther away from the center.
 As we consider individual stellar profiles of galaxies across distinct DM scenarios, a notable finding emerges.
-In the DM scenario, the stellar profiles exhibit a flatter behaviour in the inner regions, in alignment with the expected pattern of stars tracing the presence of a soliton in the core.
-In contrast, both WDM and CDM scenarios lack such a soliton, resulting in a cuspy inner profile for these scenarios.
-While the primary focus of this study lies in the comparison of evolved cosmological structures between DM and CDM, it is reasonable to anticipate that the large-scale structures in WDM would bear similarities to DM.
+In the DM scenario, the stellar profiles exhibit a flatter behavior in the inner regions, in alignment with the expected pattern of stars tracing the presence of a soliton in the core.
+In contrast, both “WDM” and CDM scenarios lack such a soliton, resulting in a cuspy inner profile.
+While the primary focus of this study lies in the comparison of evolved cosmological structures between DM and CDM, it is reasonable to anticipate that the large-scale structures in “WDM” would bear similarities to DM.
 As illustrated in Mocz et al.
-(2019), the resulting structures are remarkably akin for WDM and DM at z=5.56, despite the more pronounced differences in their internal halo structures.
+(2019), the resulting structures are remarkably akin for “WDM” and DM at z=5.56, despite the more pronounced differences in their internal halo structures.
+5.2 Stellar morphology
+In this section, we compare the stellar distribution of each Galaxy across different DM simulations.
+The main objective is to detect any potential imprint in the stellar structure of these Galaxies resulting from their respective dark matter models.
+Since the data is limited to high values of z, it is important to note that only the results for Galaxy 1 can be considered entirely reliable, as it is the only one to have a fully virialized halo.
+Figure 2 presents the stellar morphology at three projections for the isolated G1 at .
+The main panels in this plot display the stellar distribution for the entire Galaxy at a scale of approximately 55 kpc/h, while the additional subpanels provide a closer view of the central regions of the Galaxy (5 kpc/h).
+This allows for a detailed comparison between different stellar cores.
+Figure 2 distinctly illustrates variations among three DM models, revealing that stars exhibit a more asymmetric structure in DM compared to a slightly symmetric shape seen in “WDM” and progressively a more symmetric shape in CDM, aligning well with theoretical predictions (Gao & Theuns 2007; Schive et al.
+2020).
+Notably, there is a significant discernible difference between the stellar distributions in “WDM” and DM.
+Despite appearing similar in scale in Figure 1, stars in “WDM” are actually more flattened than in DM.
+Upon closer examination in the subpanels with zoomed data, stars exhibit a spherical form in the DM core, but a more elliptical shape toward the center for both CDM and “WDM”.
+The plot effectively highlights the disparities between the solitonic core (DM) and cuspy dark matter centers (CDM)/“WDM”.
+Additionally, aside from having fewer stars than CDM and “WDM”, DM has a smaller proportion of stars in the central regions (see Table 4).
+This suggests a more extended stellar profile around the soliton, in line with recent discoveries of halos around some studied dwarfs (Chiti et al.
+2021; Collins et al.
+2021).
+These extended halos are inherent to DM, composed of excited states above the ground state soliton, as evidenced by the NFW form, as predicted by the DM simulations (Schive et al.
+2014a).
+This reflects the fundamentally nonrelativistic nature of DM.
+Moving to Figure 10, we observe “WDM”–CDM data for z 2.23.
+The results are comparable for both models when compared to the data from redshift z 5.56, showing an expected symmetry.
+The additional subpanels in Figure 2 prominently display the predicted asymmetry for DM, a distinctive feature not observed in the symmetric distributions of stars in “WDM” and CDM models (Schive et al.
+2020).
+This asymmetry in DM arises from the solitonic characteristic motion, which includes a random walk caused by interference on the de Broglie scale, where the spatial position of the center of the soliton changes with time, and the inherent oscillation ’jumping’ behavior (soliton oscillations) on the scale of the de Broglie wavelength or the core radius (), owing to its standing wave nature (Li et al.
+2021; Chiang et al.
+2021).
+The jumps occur over a timescale of years (Pozo et al.
+2023) for a dSph with 0.3 kpc and 10 km/s, values closely matching those for this Galaxy (as referred in Tables 1 and 2).
+Notably, this timescale is equivalent to the stellar crossing time of the soliton, .
+In essence, this means that stars “sense” the movement of the soliton and respond accordingly, resulting in the observed asymmetry (Li et al.
+2021; Chiang et al.
+2021).
+We are exclusively addressing the observed comet tail in the core of the Galaxy.
 5.3 Core-halo structure
-Below, we explore the presence of core-halo structures across various galaxies.
+Below, we explore the presence of core-halo structures across various Galaxies. Before we proceed, it is important to define some concepts for the analysis.
+We define the core-halo structure as the characteristic resulting profile observed in all ideal DM dark matter models, featuring an inner flat core that can be fitted with a soliton and an outer characteristic asymptotic profile.
+The transition point, , is simply the radius at which this change of regime occurs.
+This characteristic profile has been observed in the stellar profiles of many Local Group galaxies (Pozo et al.
+2024, 2023), motivating the search for such structure in the stellar profiles of the simulations presented in this work.
+To analyze the compatibility of the simulated stellar profiles with the searched core-halo structure, we fit each stellar profile using a core model for the inner region (Eq.
+8 for DM, and an isothermal model (Binney & Tremaine (2008), Eq.
+9) for “WDM” and CDM, since Eq.
+8 assumes a galaxy hosting a soliton, making it incompatible with “WDM” and CDM).
+For the outer halo, we use an NFW-like profile (Navarro & Frenk 1996).
+A clear intersection between these two profiles indicates a high level of compatibility with the investigated core-halo structure.
+Since the data is limited to high values of z, it is important to note that only the results for Galaxy 1 can be considered entirely reliable, as it is the only one to have a fully virialized halo.
 5.3.1 Signature of the core-halo structure
@@
 Moving from left to right, we present the first, second, and third Galaxy, respectively, as was previously pointed out in Figure 1.
-In the first row, a notable difference is seen between the stellar profiles in the DM scenario and CDM.
+In this figure, a notable difference is seen between the stellar profiles in the DM scenario and CDM.
 This highlights how stellar profiles can vary and evolve differently in distinct DM scenarios, suggesting they could serve as a valuable tool for analyzing DM signatures in dwarf galaxies.
-The second row provides a closer look at the inner part of the galaxy and offers a clear test to assess the compatibility of the profiles with a core-halo structure.
-To this end, we plot the ideal solitonic and NFW profiles that would correspond to the inner core and the characteristic asymptotic profile, respectively.
-It is worth noting that only in the case of DM do the NFW and solitonic profiles intersect while still fitting the stellar profile.
-This is never a possibility for any CDM galaxy.
+If we take a closer look at the inner part of the Galaxy we can make a clear test to assess the compatibility of the profiles with a core-halo structure.
+To this end, we plot the ideal solitonic/isothermal and NFW profiles that would correspond to the inner core and the characteristic asymptotic profile, respectively.
+It is worth noting that only in the case of DMG1, for figure 3, do the NFW and solitonic/isothermal profiles intersect while still fitting the stellar profile.
+This is never a possibility for any CDM Galaxy, where the entire profile can be fitted by a single NFW or isothermal profiles.
 This result may indicate the potential presence of a primordial core-halo stellar structure.
@@
 This is a consequence of the anticipated cusp, whereas a soliton is required for DM to establish its flat inner profile.
-The reason why both stellar profiles appear to coincide with a flat core for galaxy one in the inner part could be attributed to the widely suggested notion that baryonic feedback may lead to the development of a core, even in the CDM scenario (Chan et al., 2015; Pontzen & Governato, 2012).
-This is more plausible at this redshift for G1 than for G2 and G3, given its greater mass as the inner slope of the DM halo profiles shows a strong mass dependence (Chan et al., 2015).
+The reason why both stellar profiles appear to coincide with a flat core for Galaxy 1 in the inner part could be attributed to the widely suggested notion that baryonic feedback may lead to the development of a core, even in the cold dark matter (CDM) scenario (Chan et al.
+2015; Pontzen & Governato 2012).
+This is more plausible at this redshift for G1 than for G2 and G3, given its greater mass, as the inner slope of the dark matter (DM) halo profiles shows a strong mass dependence (Chan et al.
+2015).
+However, CDM G2 also requires a cored fit, albeit much smaller than the large flat inner core of DM G2.
 It is likely the same reason why the core-halo structure had sufficient time to form only in this Galaxy.
@@
 Below this radius, cautious must be taken in interpreting the data.
-Next, we proceed with comparing the stellar profiles at between the WDM and CDM simulations, following the approach outlined in the previous section.
+Next, we proceed with comparing the stellar profiles at between the “WDM” and CDM simulations, following the approach outlined in the previous section.
 It is worth noting that there was no available data at this redshift for DM.
-Despite the significant differences between WDM and DM on small scales where the core-halo structure is anticipated, their similar large-scale evolution and initial conditions could make the comparison between WDM and CDM valuable for future studies.
+Despite the significant differences between “WDM” and DM on small scales where the core-halo structure is anticipated, their similar large-scale evolution and initial conditions could make the comparison between “WDM” and CDM valuable for future studies.
 The resulting extracted stellar profiles, shown in Figure 4, seem to reaffirm the same findings as the previous redshift for CDM.
 In CDM, there is no evidence or imprint of a core-halo structure identified in the halos of the three simulated Galaxies.
-In contrast, one WDM halo exhibits a potentially compatible core-halo stellar profile, particularly in a primordial context, suggesting that they may still be in the process of forming for galaxies one and two.
-Furthermore, a well-defined core-halo structure is evident for the third galaxy, characterized by a notable density drop from the core to the halo, along with the distinctive change of regime point (transition point, ) marking the boundary between the inner core and the outer halo (as depicted in the lower panels of Figure 3).
-This is observable where the pure NFW and solitonic profiles intersect.
-It is interesting to observe that for WDM, the core-halo structure emerges earlier in a lower mass galaxy, such as galaxy three, in comparison to the more massive galaxy one.
+In contrast, one “WDM” halo exhibits a potentially compatible core-halo stellar profile, particularly in a primordial context, suggesting that they may still be in the process of forming for Galaxies one and two.
+Furthermore, a well-defined core-halo structure is evident for the third Galaxy, characterized by a notable density drop from the core to the halo, along with the distinctive change of regime point (transition point, ) marking the boundary between the inner core and the outer halo.
+This is observable where the pure NFW and cored profiles intersect.
+It is interesting to observe that for “WDM”, the core-halo structure emerges earlier in a lower mass galaxy, such as Galaxy three, in comparison to the more massive Galaxy one.
 This structure should be established by a more pronounced core, likely a combination of a dense soliton and potentially baryonic feedback, both more prominent in heavier galaxies.
-Nevertheless, it is crucial to remember that WDM lacks a solitonic core.
-Indeed, the presence of a core-halo structure remains a topic for future analysis.
-5.2.2 Stellar, dark matter and gas distribution
+Nevertheless, it is crucial to remember that “WDM” lacks a solitonic core.
+The presence of a core-halo structure remains intriguing for “WDM” and that’s the main reason why we only proceed to evaluate it for DMG1, figure 6.
+5.3.2 Stellar, dark matter, and gas distribution
 Here we compare distinct distributions of stellar, gaseous, and dark matter profiles in all three different simulated Galaxies across different DM scenarios.
-The aim is to extend the former comparison made only for the stellar profiles, done in Figure 2 and 3, to further include the DM and gaseous distributions as well.
+The aim is to extend the former comparison made only for the stellar profiles, done in Figure 3 and 4, to further include the DM and gaseous distributions as well.
 Figure 5 presents these profiles at .
@@
 Notably, a common feature is the presence of a core-halo transition in the DM profile from DM and gas profiles, even if it is only visibly prominent in the case of stars in G1.
-In comparison with DM, WDM and CDM exhibit strikingly similar cuspy stellar and DM profiles, while the gas profile resembles that of the DM case.
-From the plot, it is inferred that in DM model both of the stellar and gaseous profiles demonstrate similar core radii and transition points compared with the DM profile.
+In comparison with DM, “WDM” and CDM exhibit strikingly similar cuspy DM profiles with smaller flat inner regions, while the gas profile resembles that of the DM case.
+From the plot, it is inferred that in DM model both of the stellar and gaseous profiles demonstrate similar core radii and transition points compared with the DM profile (for G1).
 This suggests that stars, rather than gas, may hold the key to determine the underlying DM model.
-Another crucial observation is the nearly identical core size and transition point in stars and gas, closely mirroring that reproduced by the DM particles.
-This implies a potential baryonic-dark matter connection and hints at the possibility of baryonic tracking of dark matter particles (Pozo et al., 2021; Pozo et al., 2022; Pozo et al., 2020, 2023).
+Another crucial observation is the nearly identical core size and transition point in stars and gas, closely mirroring that reproduced by the DM particles (for DM G1).
+This implies a potential baryonic-dark matter connection and hints at the possibility of baryonic tracking of dark matter particles (Pozo et al.
+2021, 2022, 2024, 2023).
+It is also important to remark that stellar profiles exhibit differences among the three DM models, with larger flat cores and more prominent declines in the DM cases compared to the smoother shapes of “WDM”–CDM.
+This reveals that stellar profiles evolve and behave slightly differently depending on the DM models, reinforcing their potential utility in distinguishing between the underlying DM models.
 | Combinations | ||||
 | (kpc) | (kpc) | eV | ||
-| DM | 5.56 | |||
-| WDM | 2.23 | |||
+| DMG1 | 5.56 | |||
+| dSphBoth | 0 | |||
 | dSph | 0 | |||
-| dSph | 0 | |||
-| dSph | 0 | |||
+| dSphAndromeda | 0 | |||
 | Draco | 0 | - | ||
@@
 | kpc | kpc | |||
-| DM | 8.3 | - | 52 | - |
-| DM | 5.5 | - | 45 | - |
-| DM | 5.9 | - | 46 | - |
-| CDM | 17 | 39.5 | 65 | 83 |
-| CDM | 4.5 | 26.8 | 42 | 73 |
-| CDM | 8.3 | 26.8 | 52 | 73 |
-| WDM | 13 | 35.5 | 60 | 80 |
-| WDM | 7.3 | 38.2 | 49 | 82 |
-| WDM | 6.9 | 24.8 | 48 | 70 |
+| DMG1 | 8.3 | - | 52 | - |
+| DMG2 | 5.5 | - | 45 | - |
+| DMG3 | 5.9 | - | 46 | - |
+| CDMG1 | 17 | 39.5 | 65 | 83 |
+| CDMG2 | 4.5 | 26.8 | 42 | 73 |
+| CDMG3 | 8.3 | 26.8 | 52 | 73 |
+| “WDM”G1 | 13 | 35.5 | 60 | 80 |
+| “WDM”G2 | 7.3 | 38.2 | 49 | 82 |
+| “WDM”G3 | 6.9 | 24.8 | 48 | 70 |
 5.3.3 Estimation of the core-halo parameters
-We now proceed with MCMC-based DM profile fits (Eq.
-(7)) to the DM, as shown in Figure 5 ( Corner plot Figure 14).
+We now proceed with MCMC-based DM stellar profile fits (Eq.
+(7),) to the DMG1, as shown in Figure 6 ( Corner plot Figure 13).
 The core radius (), as defined in Eq.
@@
 The transition radius (), defining the point of density transition between the soliton and NFW profiles, is the only other free parameter.
-We vary this within a prior range determined by DM simulations, specifically to times (Schive et al., 2014a, b).
-The boson mass stands as the sole fixed parameter, maintaining a value of eV, consistent with that used for the simulations (Mocz et al., 2020).
+We vary this within a prior range determined by DM simulations, specifically to times (Schive et al.
+2014a, b).
+The boson mass stands as the sole fixed parameter, maintaining a value of eV, consistent with that used for the simulations (Mocz et al.
+2020).
 The best-fit yields a resulting profile with a total mass of , aligning well with the computed mass from Mocz et al.
@@
 It is worth noting that while it may not match the accuracy of the green profile with a freely varying mass, it still demonstrates excellent agreement.
-Furthermore, the resulting values for kpc and kpc, fall within the expected range for such dSphs, and are very close to the directly inferred soliton radius from the simulation ( kpc (Pozo et al., 2020, 2023)).
-This concordance between the suggested soliton size based on stellar behaviour and the actual size of the simulated DM halo soliton may hold significant implications for our understanding of how stellar behaviour is influenced by dark matter.
-Next, we proceed with fitting a stellar core-halo profile for another case where a core-halo shape is evident: the WDM G3 at .
-It is crucial to note that in principle fitting a core-halo profile to WDM, without a physical soliton, might be challenging.
-In this sense, we are careful not to directly link the core-halo structure to a soliton but rather to an effective profile with a transition from the core to the halo in WDM at this lower redshift.
-The resulting values for core radius () as well as the transition point () are summarized in Table 1, alongside other pertinent variables, with the resulting profiles depicted in Figure 6 ( Corner plot Figure 15).
+Furthermore, the resulting values for kpc and kpc, fall within the expected range for such dSphs, and are very close to the directly inferred soliton radius from the simulation ( kpc (Pozo et al.
+2024, 2023)).
+This concordance between the suggested soliton size based on stellar behavior and the actual size of the simulated DM halo soliton may hold significant implications for our understanding of how stellar behavior is influenced by dark matter.
+The resulting values for core radius () as well as the transition point () are summarized in Table 1, alongside other pertinent variables.
 It is noteworthy that these parameter values fall within the expected range when compared to those extracted from observed profiles of Local Group classical dSphs.
 This is particularly striking considering they were analyzed at different redshifts and are still in the process of the formation.
-Table 1 also provides the parameter values for the two extracted galaxies and the observed dSphs in the Local Group (Pozo et al., 2023).
+Table 1 also provides the parameter values for the two extracted galaxies and the observed dSphs in the Local Group (Pozo et al.
+2023).
 The core radius values are in excellent agreement with the mean values obtained for such dSphs in the Local Group.
 The isolated Galaxy (G1) from the DM simulation data (, Table 1) exhibits a core radius akin to that of two Milky Way satellite galaxies, Leo II and Sculptor.
-Conversely, the third galaxy from the WDM simulation data (WDM, Table 1) aligns exceptionally well with the observed cores of Sculptor and And IX.
+Conversely, the third Galaxy from the “WDM” simulation data (third panel of Figure 4) seems to align well with the observed cores of Sculptor and And IX.
 Nonetheless, there are slight disparities in the transition point, which is reasonable given the difference in redshift and the fact that these simulated Galaxies are still in the process of formation.
-The observed increase in between and WDM could be attributed to the lower redshift of WDM, which aligns with the expectation of a larger .
 Notably, despite the significant redshift difference (almost 6 between the observed galaxies and the observable Local Group dwarfs), the values remain similar.
@@
 5.3.4 Comparison with observations
-We proceed by comparing the extracted profile with two known dwarf galaxies, Cetus and Andromeda XXI, that orbit the Milky Way and Andromeda, respectively, and exhibit the most similar density gaps (See Figure 7).
+We proceed by comparing the extracted DM and CDM G1 profiles with two known dwarf galaxies, Cetus and Andromeda XXI, that orbit the Milky Way and Andromeda, respectively, and exhibit the most similar density gaps (see Figure 7).
 In the left panel, the profiles are normalized based on their stellar peak density, while the right panel is additionally normalized by their respective transition radii.
-The black profiles represent the mean stellar profile of all Local Group dwarf spheroidal galaxies (Pozo et al., 2023).
-The right panel demonstrates that the extracted DM stellar profile (blue profile) aligns well with a real observed stellar profile, Andromeda XXI (red profile).
-This indicates that the extracted profile can reproduce observed dSph structures, in contrast to the CDM profiles.
+The black profiles represent the mean stellar profile of all Local Group dwarf spheroidal galaxies (Pozo et al.
+2023).
+The right panel demonstrates that the extracted DM stellar profile (blue profile) aligns well with a real observed stellar profile, Andromeda XXI (orange profile).
+This indicates that the extracted profile can reproduce observed dSph structures, in contrast to the CDM profile, here represented in red.
 Furthermore, this panel reveals that the inner portions of the profiles are in good agreement, suggesting that they are consistent with a solitonic core.
-However, it’s crucial to note that Andromeda XXI and the extracted profile are at different redshifts, 5.5 and 0 respectively.
+However, it is crucial to note that Andromeda XXI and the extracted profile are at different redshifts, 5.5 and 0 respectively.
 This indicates that the extracted profile still has much more time to evolve, and its final core-halo structure at redshift 0 could substantially change over time.
-| Halos | (-2<[Fe/H]<0) | (-4<[Fe/H]<-2) | (-6<[Fe/H]<-4) | (-8<[Fe/H]<-6) | |
-| DM | 12340 | 93.12 | 4.91 | 0.71 | 1.26 |
-| DM | 1950 | 80.62 | 14.82 | 1.44 | 3.12 |
-| DM | 1550 | 77.22 | 16.91 | 1.68 | 4.19 |
-| CDM | 21206 | 95.62 | 3.61 | 0.36 | 0.41 |
-| CDM | 7572 | 88.55 | 9.17 | 1.12 | 1.16 |
-| CDM | 3417 | 83.79 | 12.38 | 1.82 | 2.01 |
-| WDM | 28328 | 89.48 | 8.43 | 0.76 | 1.33 |
-| WDM | 3793 | 79.25 | 14.79 | 2.14 | 3.82 |
-| WDM | 5122 | 78.21 | 16.34 | 1.44 | 4.01 |
-| Halos | ||||||||
-| DM | 12340 | 11.57 | 8.02 | 0.92 | 48.72 | 37.41 | 10.46 | 3.41 |
-| DM | 1950 | 5.66 | 2.82 | 0.16 | 72.10 | 16.41 | 7.23 | 4.26 |
-| DM | 1550 | 5.10 | 2.89 | 0.15 | 64.45 | 27.36 | 8.19 | 0 |
-| CDM | 21206 | 13.21 | 5.58 | 0.74 | 48.60 | 37.27 | 7.75 | 6.38 |
-| CDM | 7572 | 24.57 | 8.24 | 2.02 | 29.32 | 30.78 | 14.05 | 25.85 |
-| CDM | 3417 | 7.99 | 10.87 | 0.87 | 35.91 | 24.00 | 9.51 | 30.58 |
-| WDM | 28328 | 24.56 | 13.75 | 3.37 | 38.04 | 36.55 | 12.05 | 13.36 |
-| WDM | 3793 | 10.18 | 6.14 | 0.63 | 63.88 | 20.43 | 6.52 | 9.17 |
-| WDM | 5122 | 16.08 | 9.31 | 1.50 | 49.57 | 38.54 | 11.21 | 0.68 |
-5.3 Stellar Morphology
-In this section, we will compare the stellar distribution of each galaxy across different DM simulations.
-Figure 8 presents the stellar morphology at three projections for the isolated G1 at .
-The main panels in this plot display the stellar distribution for the entire galaxy at a scale of approximately 55 kpc/h, while the additional sub-panels provide a closer view of the central regions of the galaxy (5 kpc/h).
-This allows for a detailed comparison between different stellar cores.
-Figure 8 distinctly illustrates variations among three DM models, revealing that stars exhibit a more asymmetric structure in DM compared to a slightly symmetric shape seen in WDM and progressively a more symmetric shape in CDM, aligning well with theoretical predictions (Gao & Theuns, 2007; Schive et al., 2020).
-Notably, there is a significant discernible difference between the stellar distributions in WDM and DM.
-Despite appearing similar in scale in Figure 1, stars in WDM are actually more flattened than in DM.
-Upon closer examination in the subpanels with zoomed data, stars exhibit a spherical form in the DM core, but a more elliptical shape towards the center for both CDM and WDM.
-The plot effectively highlights the disparities between the core and cuspy centers, indicating that stars tend to concentrate in smaller cores in DM model, contrasting the cusps observed in CDM/WDM.
-Additionally, aside from having fewer stars than CDM and WDM, DM has a smaller proportion of stars in the central regions (see Table 4).
-This suggests a more extended stellar profile around the soliton, in line with recent discoveries of halos around some studied dwarfs (Chiti et al., 2021; Collins et al., 2021).
-These extended halos are inherent to DM, composed of excited states above the ground state soliton, as evidenced by the NFW form, as predicted by the DM simulations (Schive et al., 2014a).
-This reflects the fundamentally non-relativistic nature of DM.
-Moving to Figure 11, we observe WDM/CDM data for z 2.23.
-The results are comparable for both models when compared to the data from redshift z 5.56, showing an expected symmetry.
-The additional sub-panels in Figure 8 prominently display the predicted asymmetry for DM, a distinctive feature not observed in the symmetric distributions of stars in WDM and CDM models (Schive et al., 2020).
-This asymmetry in DM arises from the solitonic characteristic ‘jumping’ behaviour (soliton oscillations) on the scale of the de Broglie wavelength , or the core radius (), owing to its standing wave nature (Li et al., 2021; Chiang et al., 2021).
-The jumps occur over a timescale of years for a dSph with 0.3 kpc and 10 km/s, values closely matching those for this galaxy (as referred in Tables 1 and 2 and check out Pozo et al.
-(2023) for a deeper explanation of ).
-Notably, this timescale is equivalent to the stellar crossing time of the soliton, .
-In essence, this means that stars ‘sense’ the movement of the soliton and respond accordingly, resulting in the observed asymmetry (Li et al., 2021; Chiang et al., 2021).
-5.3.1 Stellar age vs metallicity
+5.4 Stellar age vs. metallicity
 Metallicity plays a crucial role in unraveling the galaxy history.
 Here, we explore the relationship between the stellar metallicity and age.
-The corresponding data is displayed in the right columns of Figures 9 and 12.
+The corresponding data is displayed in the right columns of Figures 8 and 11.
 These figures illustrate the [Fe/H] versus stellar age profiles for G1 at the and for Galaxies 1 and 3 at the .
 We chose to focus on these specific Galaxies to emphasize the most pertinent data associated with instances where the potential presence of a core-halo structure has been identified.
-We observe that in DM and WDM, the stellar metallicity exhibits a smoother decline as the stellar age increases.
-On the other hand, CDM displays a more pronounced breaks and a more discrete age vs.
+We observe that DM and “WDM” have pretty similar profiles with the metal-poor concentrations at similar ages, relaying a similar stellar history.
+On the other hand, CDM displays a more pronounced breaks(more discrete) age vs.
 [Fe/H] profile, suggesting that its associated stars may have diverse origins.
 This disparity could arise from a combination of a few different factors.
-In CDM, the stellar cycle is more extended, with the first stars forming around z30 and continuing up to 9.8 billion years (as seen in Figure 9).
-In contrast, in WDM/DM, star formation begins around z13 (Mocz et al., 2019, 2020).
-Additionally, DM structures form earlier in CDM compared to WDM/DM, a consequence of the small-scale cut-off in the power spectrum.
-Continuing with this idea, all three galaxies exhibit a peak of star formation around 9.6 billion years and 9.4 billion years, with the former being the predominant one for WDM/DM and the latter for CDM.
-Additionally, CDM is the only model with an additional small peak at 9.5 billion years.
-It is evident that all three models display a moderate to strong negative correlation between stellar age and metallicity.
-Similar trends and predictions are confirmed in Figure 12 at .
-The observations in Figures 9 and 12 suggest that metal-rich stars tend to be more concentrated in the central regions of the galaxies, while metal-poor stars are located farther out, which is consistent with previous studies (Dietz et al., 2020; Santucci & otehrs, 2020).
-Additionally, the majority of stars are concentrated within distances compatible with the core-halo structure (r < 2 kpc, marked with a vertical dashed line).
-Of particular interest is that the galaxies where a compatible core-halo structure has been identified (DM at and WDM at ) exhibit a radially-offset region of stellar-density enhancement, which is especially pronounced in WDM at (as seen in Figure 12).
-Notably, WDM is the only case where this concentration of stars extends beyond the theoretical limit of the core-halo structure (r > 2 kpc).
+In CDM, the stellar cycle is more extended, with the first stars forming around z30 ( stellar age 0.2) and continuing up to 1.2 billion years (as seen in Figure 8).
+In contrast, in “WDM”/DM, star formation begins around z13 ( stellar age 0.5) (Mocz et al.
+2019, 2020).
+Additionally, DM structures form earlier in CDM compared to “WDM”/DM, a consequence of the small-scale cutoff in the power spectrum.
+Continuing with this idea, all three Galaxies exhibit a peak of star formation around 0.6 billion years and 1.0 billion years, with the former being the predominant one for “WDM”/DM and the latter for CDM.
+Additionally, CDM is the only model with an additional small peak at 0.85 billion years.
+As the current data does not cover more than 1 billion years, we cannot confirm the expected negative correlation between stellar age and metallicity.
+Similar data and predictions are shown in Figure 11 at .
+The observations in Figures 8 and 11 suggest that metal-rich stars tend to be more concentrated in the central regions of the Galaxies, while metal-poor stars are located farther out, which is consistent with previous studies (Dietz et al.
+2020; Santucci & otehrs 2020).
+Additionally, the majority of stars are concentrated within distances compatible with the core-halo structure (r ¡ 2 kpc, marked with a vertical dashed line).
+Of particular interest is that the Galaxies where a compatible core-halo structure has been identified (DMG1 at and “WDM”G3 at ) exhibit a radially offset region of stellar-density enhancement, which is especially pronounced in “WDM”G3 at (as seen in Figure 11).
+Notably, “WDM”G3 is the only case where this concentration of stars extends beyond the theoretical limit of the core-halo structure (r ¿ 2 kpc).
 This peculiarity may be attributed to the filamentary environment observed for this Galaxy, an environment that cannot be reproduced in the CDM model (as shown in Figure 1).
@@
 In summary, galaxies with a higher concentration of stars extending along their halos, rather than being centrally concentrated, seem to be more likely to develop the core-halo structure.
-5.3.2 Spatial distribution of metallicity and age
-Figures 10 and 13 investigate the correlation between the mean stellar metallicity and the stellar radii for galaxies from different DM models at both redshifts.
+Figures 9 and 12 investigate the correlation between the mean stellar metallicity and the stellar radii for Galaxies from different DM models at both redshifts.
 The mean stellar age is depicted in the color bars.
-The analysis of the correlation between stellar age and the spatial distribution of metallicity (as shown in Figures 10 and 13) reveals distinct differences among the three DM models.
-One notable distinction is that in CDM and WDM stars are relatively more metal-rich ([Fe/H] -1.0) compared to DM ([Fe/H] -1.25).
-Furthermore, CDM and WDM exhibit a smoother, continuous decline in metallicity from the center to the outer halo, whereas DM displays a more discrete profile, with noticeable drops in metallicity around 0.3 kpc and 0.5 kpc (as observed in the right panel of the first row in Figure 10).
-Regarding the spatial distribution of stellar age, WDM exhibits older stars concentrated in the center, while the mean age decreases towards the outer parts of the halo.
+The analysis of the correlation between stellar age and the spatial distribution of metallicity (as shown in Figures 9 and 12) reveals distinct differences among the three DM models.
+One notable distinction is that in CDM and “WDM” stars are relatively more metal-rich ([Fe/H] -1.0) compared to DM ([Fe/H] -1.25).
+Furthermore, CDM and “WDM” exhibit a smoother, continuous decline in metallicity from the center to the outer halo, whereas DM displays a more discrete profile, with noticeable drops in metallicity around 0.3 kpc and 0.5 kpc (as observed in the right panel of the first row in Figure 9).
+Regarding the spatial distribution of stellar age, “WDM” exhibits older stars concentrated in the center, while the mean age decreases toward the outer parts of the halo.
 This trend is opposite to what is observed for CDM and DM.
-However, it is important to note that for DM, the mean age is lower compared to the mean age observed for CDM, which aligns with the predicted stellar history between both models (Mocz et al., 2019, 2020).
+However, it is important to note that for DM, the mean age is lower compared to the mean age observed for CDM, which aligns with the predicted stellar history between both models (Mocz et al.
+2019, 2020).
 Tables 3 and 4 provide additional data regarding the percentage of metallicity and age in the stellar profiles of all the Galaxies.
@@
 6 Discussion and conclusions
-The comparison between the core-halo structures of WDM and DM simulations with the CDM predictions reveals intriguing distinctions.
-In the case of WDM and DM, the core-halo structure appears to naturally stem from their respective power spectrum small-scale cut-off, whereas for CDM this structure is not anticipated.
-This notion could be further examined by tracking the stars of each galaxy across different epochs and observing their initial positions to scrutinize in more detail whether the stars appearing in the outer extended halos were born within their halo or in the outer filamentary structures.
-This observation, combined with other disparities between CDM predictions and observations, has spurred the investigation of alternative DM candidates like WDM and DM.
-Moreover, the fact that the stellar profiles in each DM scenario adhere to the predicted behaviours, with a cuspy profile for WDM and CDM, and a flat core structure for DM, supports the notion that stars trace DM.
-Figure 4 vividly illustrates that the DM, stellar, and the gaseous density profiles (especially stellar and DM) differ significantly in each DM scenario, and that the resulting stellar profile varies depending on the DM profile (cuspy/core discrepancy and core-halo structure).
+The comparison between the stellar core-halo structures (not the stellar core itself, but the observed inner core and outer halo structure in Figures 3 and 6, first detected in Pozo et al.
+(2024)) of DM simulations with the CDM predictions reveals intriguing distinctions.
+In the case of DM, the core-halo structure appears to naturally stem from the also observable DM core-halo structure (see the first panel of the top row of figure 5), whereas for CDM, this structure is not anticipated.
+Moreover, as seen in figure 5, it is clear that in DM the N-body simulations exhibit a cuspy DM profile, whereas only DM presents solitonic cores.
+Any core formation in CDM and “WDM” seems to be driven by baryonic physics rather than being directly related to the DM component.
+Therefore, baryonic physics is the most likely explanation for the transformation of the observed DM cusp in “WDM”–CDM into stellar cores.
+This notion could be further examined by tracking the stars of each galaxy across different epochs and observing their initial positions to scrutinize in greater detail whether the stars appearing in the outer extended halos were born within their halo or in the outer filamentary structures.
+This observation, combined with other disparities between CDM predictions and observations, has spurred the investigation of alternative DM candidates such as “WDM” and DM.
+Moreover, the fact that the stellar profiles in each DM scenario follow the predicted behaviors, with a cuspy profile for CDM and a flat or prominent core structure for DM, supports the notion that stars could trace DM.
+Figure 5 vividly illustrates significant differences in the density profiles of DM, stars, and gas in each DM scenario.
+The resulting stellar profile varies depending on the DM model, indicating a cuspy–core discrepancy and core-halo structure.
+Finally, this work aims to clarify how the core-halo structure can evolve and how definitive it can be to support any dark matter model.
+It can be observed that the core-halo structure appears in the DM dark matter density profiles but never for CDM–“WDM”.
+This point alone might not be significant enough if it were not for the fact that such core-halo structures, which we are observing for dark matter, also appear in the DM Galaxy 1 stellar profile and probably in Galaxy 3 for “WDM” at z=2.23, as can be seen in Figures 3 and 4.
+However, it never appears for CDM.
+This strongly supports the assertion that the stellar profiles show discrepancies for the three different dark matter models, as we claim in this work.
 This underscores the potential significance of studying the properties of stars to achieve a deeper comprehension of the characteristics of DM.
-Nonetheless, it is also true that in CDM and WDM both of the stellar and DM profiles appear to shift from a cuspy profile to cored one when decreasing the redshift (though never becoming as flat as DM predictions, which exhibit a better consistency with the observations (Sanchez Almeida et al., 2023)).
-This transformation from cuspy to cored profiles for CDM/WDM seems to align with the asserted impact of baryonic feedback in generating cores from originally cuspy halos.
-Feedback induces bursts of star formation and outflows, modifying the DM distribution and mitigating small-scale issues in CDM (Read & Gilmore, 2005; Amorisco et al., 2013; Chan et al., 2015), potentially resolving the cuspy-core problem.
+Nonetheless, it is also true that in CDM and “WDM” the stellar and DM profiles both appear to shift from a cuspy profile to a cored one when decreasing the redshift, although they never become as flat and prominent as the DM predictions, which exhibit a better consistency with the observations (Sanchez Almeida et al.
+2023).
+This transformation from cuspy to cored profiles for CDM–“WDM” seems to align with the asserted impact of baryonic feedback in generating cores from originally cuspy halos.
+While in the case of “WDM”, the inner cores are claimed as well (Macciò et al.
+2012; Lovell et al.
+2014; Bode et al.
+2001; Destri et al.
+2013).
+Feedback induces bursts of star formation and outflows, modifying the DM distribution and mitigating small-scale issues in CDM (Read & Gilmore 2005; Amorisco et al.
+2013; Chan et al.
+2015), potentially resolving the cuspy-core problem.
 Furthermore, the simulations conducted by Mocz et al.
-(2019, 2020) were highly informative, demonstrating that baryonic feedback may not possess sufficient strength to convert cusps into cores for z>6, but it does open up possibilities for lower redshifts.
-This coincides with our findings, where the cusps observed for CDM/WDM at z5.56 seem to be alleviated around z2.23.
-Certainly, the detection of the core-halo structure in WDM simulations, despite the absence of a soliton, is a topic of discussion.
-It seems that the smoothness of the cut-off in the initial dark matter power spectrum, which is likely more pronounced for WDM, may play a role in the formation of the core-halo structure in the stellar profiles.
-This suggests that the soliton itself may not be the sole factor determining the presence of the core-halo structure.
-However, it is worth noting that the presence of a soliton can be important as the transition feature in the DM and star profiles is a result of the power spectrum cut-off.
-The presence of a soliton may have a more notable impact on the early formation of the core-halo structure in DM (at ) compared to WDM (at ).
-Indeed, the small-scale cut-off of the initial power spectrum appears to be the pivotal factor in determining the core-halo structure, as it leaves its mark on both the DM and the stellar profiles.
-The resemblance of the large-scale structures between the WDM and DM, despite their distinct physics on smaller scales, lends support to this notion.
-It is worth noting that the underlying cause of the cut-off differs for WDM and DM, with WDM exhibiting a thermal free-streaming cut-off, while DM featuring a quantum mechanical cut-off.
-However, this distinction does not seem to significantly impact larger scales where the cut-off length scale is imprinted on the profiles.
+(2019, 2020) were highly informative, demonstrating that baryonic feedback may not possess sufficient strength to convert cusps into cores for z¿6, but it does open up possibilities for lower redshifts.
+This coincides with our findings where the cusps observed for CDM–“WDM” at z5.56 seem to be alleviated around z2.23.
+Certainly, the possible detection of the stellar core-halo structure in “WDM” simulations, despite the absence of a soliton, is a topic of discussion.
+It seems that the smoothness of the cutoff in the initial dark matter power spectrum, which is likely more pronounced for “WDM”, may play a role in the formation of the core-halo structure in the stellar profiles.
+This suggests that the soliton-interference pattern change of regime itself may not be the sole factor determining the presence of the core-halo structure.
+Nevertheless, it is extremely important to note that this conclusion is a product of our poor understanding of why the stellar core-halo structure appears for “WDM” (at z = 2.23, Figure 4 Galaxy 3) as we thought that it could only appear as a consequence of the expected profile observed in all ideal DM dark matter models, featuring an inner flat core fitted with a soliton and an outer characteristic asymptotic profile.
+Even more analysis is thus required for this topic, and we conclude that it has to be a coincidence due to the absence of such structure for “WDM” dark matter density profiles (see figure 5).
+It is worth noting that the underlying cause of the cutoff differs for “WDM” and DM; “WDM” exhibits a thermal free-streaming cutoff, while DM features a quantum mechanical cutoff.
+However, this distinction does not seem to significantly impact larger scales where the cutoff length scale is imprinted on the profiles.
 The simulation data presented by Mocz et al.
-(2020) demonstrates that both of the WDM and DM yield core-halo stellar profiles, which appear to be at odds with CDM predictions.
-Further analysis is warranted to validate this finding and to probe the hypothesis that the core-halo structure arises from the small-scale cut-off of the initial power spectrum.
-On the other hand, the discrepancies found in the stellar morphology, age, and metallicity suggest the notion of a less concentrated and non-cuspy inner halo as a prerequisite for the observed core-halo structure in local dwarfs.
-This idea gains further support when we observe the higher concentration of stars in CDM/WDM models compared to DM, which exhibit a lower proportion of stars in their halo centers, indicating more extended stellar profiles.
+(2020) demonstrates that only the DM core-halo stellar profiles are reliable, due to the presence of this structure in both gas and DM, which is only noticeable in DM simulations.
+On the other hand, the discrepancies found in the stellar morphology, age, and metallicity suggest the notion of a less concentrated and noncuspy inner halo as a prerequisite for the observed core-halo structure in local dwarfs.
+This idea gains further support when we observe the higher concentration of stars in CDM–“WDM” models compared to DM, which exhibit a lower proportion of stars in their halo centers, indicating that the stellar profiles are more extended.
 Furthermore, the spatial distribution of metallicity and age indicates that the core-halo structure emerges in galaxies where the greater concentration of stars is not extremely close to the center, a condition that appears only feasible in a filamentary environment, which is challenging to replicate for CDM.
-Finally, the solitonic stellar imprint is substantiated by the observed asymmetry in the core of G1 in the DM scenario, as a consequence of the previously analyzed solitonic random walk (Li et al., 2021; Chiang et al., 2021), where stars are thought to “sense” the movement of the soliton and respond accordingly, resulting in the observed asymmetry, in stark contrast to CDM/WDM outcomes.
+Finally, the solitonic stellar imprint is substantiated by the observed asymmetry in the core of G1 in the DM scenario, as a consequence of the previously analyzed solitonic random walk (Li et al.
+2021; Chiang et al.
+2021), where stars are thought to sense the movement of the soliton and respond accordingly, resulting in the observed asymmetry, in stark contrast to CDM–“WDM” outcomes.
 This further strengthens the idea of distinct stellar behaviors for each DM model.
 Quantifying this effect with the deepest JWST images of different dwarfs at various redshifts could yield valuable insights.
-Additionally, a more detailed analysis of this soliton ‘jumping’ phenomenon could be carried out with data from lower redshifts, an area we plan to explore in future studies by tracking the motion of stars around the soliton across different redshifts.
-We conclude that CDM is under significant tension with core-halo observations.
-Moreover, there are several compelling reasons to favor DM over WDM/CDM.
-Firstly, as highlighted previously, DM seems to align better with observations in terms of the presence and structure of core-halo profiles in galaxy halos, as opposed to CDM.
+Additionally, a more detailed analysis of this soliton jumping phenomenon could be carried out with data from lower redshifts, an area we plan to explore in future studies by tracking the motion of stars around the soliton across different redshifts.
+The light boson used here, 2.5 eV, is consistent with the canonical value based the internal properties of the classical dwarf spheroidal galaxies (Schive et al.
+2014a; Chen et al.
+2017; Pozo et al.
+2023), which has been claimed to be significantly lighter than some independent constraints using the Ly-forest of eV (Iršič et al.
+2017), but it should be noted that the forest estimates are not direct and rely on analogy with “WDM” simulations as they have similar power spectrum suppressions at high frequency (from the de Broglie scale and the free streaming scale, respectively); however, the DM interference introduces additional power on small scales that is pervasive, including the low-density regions probed by the Ly-forest, where a wide range of the de Broglie scale is seen in the simulations extending to low momentum in the voids (see Figure S2 of Schive et al.
+(2014a)).
+Furthermore, the claimed discrepancy with DM is below lines widths of and only at z5, in the steeply declining Jeans turnover where the gas temperature is degenerates with the DM power spectrum, and where spatial temperature variations from AGN and cluster galaxy gas heating sources (Doughty et al.
+2025) adds to the forest variance, with widespread metal injection and wide forest gaps (Zhu.
+et al.
+2022), thus raising the observed power spectrum; however, all this is too uncertain to confidently model in forest simulations.
+Hence, reliance on Ly-forest regarding the boson mass of Wave DM is far less direct than our lensing-based estimate (Amruth et al.
+2023; Broadhurst et al.
+2024).
+However, insisting on smaller scale power can be simply achieved with an additional heavier boson species (Rogers & Peiris 2021), as in the generic string axiverse (Arvanitaki et al.
+2010).
+An additional heaver boson is also motivated by the compact scale of the new UFG class of dwarf galaxy where a boson mass for which a boson mass of eV is indicated dynamically (Pozo et al.
+2023), forming a minority of the DM, most of which is comprised of the canonical eV boson species.
+Galaxy formation in this two-boson model has been shown to be viable, capable of forming UFG dwarfs (Nhan Luu et al.
+2024).
+A lensing based estimate of eV is claimed by Powell et al.
+(2023), who argue that their reconstructed source is fully resolved with VLBI, so that sufficiently light DM would cause disruption.
+The breadth of their predicted band of DM critical curves is wide, about 25% of the Einstein radius, whereas a band narrower by a factor of 3 is predicted by Eq.
+3, and by comparable lens simulations of DM (Chan et al.
+2020b; Amruth et al.
+2023), which would leave the source unscathed, and hence insensitive to the lighter boson mass we estimate here.
+A tight constraint of ¿3eV has been claimed dynamically by Dalal & Kravtsov (2022) using two dwarf galaxies Segue I and II.
+This claim rests mainly on Segue II; the significance of their boson mass estimate relies mainly on the viability of Segue II, because its velocity dispersion is too small to be detected, unlike the 5.4 Km/s dispersion established for Segue I.
+This reliance on Segue II is evident in Figure 4 of Dalal & Kravtsov (2022).
+However, this dwarf Segue II is widely recognized as an extreme tidal remnant in the careful work of Kirby et al.
+2013, and follow-up work where its heavy progenitor origin is reinforced by star orbits(Simon 2019).
+Kirby (Kirby et al.
+2013) point out that Segue II is an extreme outlier in terms of its high metallicity and low (undetected) velocity dispersion, which they conclude implies a much more massive origin similar to the regular dSph galaxy, and also conclude that Segue II may be the barest remnant of a tidally stripped Ursa Minor-sized galaxy.
+If this is the case, it is the best example of an ultra-faint dwarf galaxy that became ultra-faint through tidal stripping.
+Thus, it is widely believed that Segue II is not representative of the class of low-mass galaxies that Dalal & Kravtsov (2022) would like to use to constrain DM.
+On the other hand, Segue I appears to be a bona fide representative of the ultra-faint and dark matter-dominated dwarf with a well-detected velocity dispersion.
+A Jeans analysis of this dwarf finds a boson mass of eV (Pozo et al.
+2023).
+This is an order of magnitude difference, and the UFD class as a whole indeed fits well with this higher boson mass.
+A case for two boson species has been made to encompass both UFD and dSPh classes by some of us (Pozo et al.
+2023).
+The argument of Dalal & Kravtsov (2022) rests on simulations that seem to underpredict the observed velocity dispersion of 3.5 km/s, so that it is not clear if Segue II alone challenges DM.
+Instead, independent Jeans analysis shows that it is consistent in terms of boson mass with the other UFD galaxies with eV (Pozo et al.
+2023) and distinct from the eV dominant light boson indicated by the dark matter dominated dSph class (Pozo et al.
+2023).
+Two such light boson species are arguably well motivated by the axiverse scenario where a discrete mass spectrum of light bosons is generic to string theory Arvanitaki et al.
+(2010), and is the subject of recent multi-boson species simulations Nhan Luu et al.
+(2024); Chan et al.
+(2020b).
+We conclude that CDM is under relative tension with stellar core-halo observations (Pozo et al.
+2024, 2023), as it is never detected for any simulated CDM halo.
+More analysis in different simulations needs to be done in the future to truly support the idea that CDM does not produce these stellar core-halo structures in galaxies.
+Moreover, there are several compelling reasons to favor DM over “WDM”–CDM.
+First, as highlighted previously, DM seems to align better with observations in terms of the presence and structure of core-halo profiles in galaxy halos, as opposed to CDM.
 Additionally, DM predicts a granularity in the lensed halos Amruth et al.
-(2023) that has not been observed in WDM simulations.
-Furthermore, the lack of detection of WDM particles in laboratory searches is a substantial drawback for this model.
+(2023) that has not been observed in “WDM” simulations.
+Furthermore, the lack of detection of “WDM” particles in laboratory searches is a substantial drawback for this model.
 In contrast, DM posits light bosons that would necessitate a much larger laboratory scale of 1 kpc for detection, making it challenging to rule out this possibility based solely on laboratory experiments.
-Overall, the combination of better agreement with observations and the lack of definitive evidence against it makes DM a promising candidate for dark matter, and a more viable alternative to WDM/CDM.
-Acknowledgements
+Overall, the combination of better agreement with observations and the lack of definitive evidence against it makes DM a promising candidate for dark matter, and a more viable alternative to “WDM”–CDM.
+Acknowledgements.
 We warmly acknowledge Douglas Finkbeiner for very fruitful conversations.
 A.P.
-is grateful for the continued support of the DIPC graduate student program as well as the Center for Astrophysics | Harvard & Smithsonian for their warm hospitality at the beginning of this project.
+is grateful for the continued support of the DIPC graduate student program as well as the Center for Astrophysics — Harvard & Smithsonian for their warm hospitality at the beginning of this project.
 R.E.
@@
 acknowledges this work was in part performed under the auspices of the U.S.
-Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, Lawrence Livermore National Security, LLC.
-Data Availability
+Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, Lawrence Livermore National Security, LLC.Data availability
 The data underlying this article will be shared on reasonable request to the corresponding author.
@@
 It is valuable to explore how these parameters interact and correlate at different redshifts, providing a more comprehensive view of the underlying dynamics.
+Figures 10, 11 and 12.
+Appendix B Tables with metallicity, age and spatial distribution of stars
+Here we present a summary in two tables ( Tables 3 and 4) of the metallicities, ages, and spatial distribution of the stars in the three simulated halos.
+These tables are interesting for quantitatively observing the distribution of stars in terms of age, metallicity, and spatial location within both the cores and halos of the simulated Galaxies across the three dark matter models.
+| Halos | (-2¡[Fe/H]¡0) | (-4¡[Fe/H]¡-2) | (-6¡[Fe/H]¡-4) | (-8¡[Fe/H]¡-6) | |
+| DMG1 | 12340 | 93.12 | 4.91 | 0.71 | 1.26 |
+| DMG2 | 1950 | 80.62 | 14.82 | 1.44 | 3.12 |
+| DMG3 | 1550 | 77.22 | 16.91 | 1.68 | 4.19 |
+| CDMG1 | 21206 | 95.62 | 3.61 | 0.36 | 0.41 |
+| CDMG2 | 7572 | 88.55 | 9.17 | 1.12 | 1.16 |
+| CDMG3 | 3417 | 83.79 | 12.38 | 1.82 | 2.01 |
+| “WDM”G1 | 28328 | 89.48 | 8.43 | 0.76 | 1.33 |
+| “WDM”G2 | 3793 | 79.25 | 14.79 | 2.14 | 3.82 |
+| “WDM”G3 | 5122 | 78.21 | 16.34 | 1.44 | 4.01 |
+| Halos | ||||||||
+| DMG1 | 12340 | 11.57 | 8.02 | 0.92 | 48.72 | 37.41 | 10.46 | 3.41 |
+| DMG1 | 1950 | 5.66 | 2.82 | 0.16 | 72.10 | 16.41 | 7.23 | 4.26 |
+| DMG1 | 1550 | 5.10 | 2.89 | 0.15 | 64.45 | 27.36 | 8.19 | 0 |
+| CDMG1 | 21206 | 13.21 | 5.58 | 0.74 | 48.60 | 37.27 | 7.75 | 6.38 |
+| CDMG2 | 7572 | 24.57 | 8.24 | 2.02 | 29.32 | 30.78 | 14.05 | 25.85 |
+| CDMG3 | 3417 | 7.99 | 10.87 | 0.87 | 35.91 | 24.00 | 9.51 | 30.58 |
+| “WDM”G1 | 28328 | 24.56 | 13.75 | 3.37 | 38.04 | 36.55 | 12.05 | 13.36 |
+| “WDM”G2 | 3793 | 10.18 | 6.14 | 0.63 | 63.88 | 20.43 | 6.52 | 9.17 |
+| “WDM”G3 | 5122 | 16.08 | 9.31 | 1.50 | 49.57 | 38.54 | 11.21 | 0.68 |
+Appendix C Corner Plots from MCMC fit
+Here we present the resulting corner plot of the performed MCMC fit corresponding to Figure 6. Figure 13.